### v2

# GD doesn’t make the cut: Three ways that non-differentiability affects neural network training

Abstract
This paper investigates the distinctions between gradient methods applied to non-differentiable functions (NGDMs) and classical gradient descents (GDs) designed for differentiable functions. First, we demonstrate significant differences in the convergence properties of NGDMs compared to GDs, challenging the applicability of the extensive neural network convergence literature based on $L-smoothness$ to non-smooth neural networks. Next, we demonstrate the paradoxical nature of NGDM solutions for $L_{1}$-regularized problems, showing that increasing the regularization penalty leads to an increase in the $L_{1}$ norm of optimal solutions in NGDMs. Consequently, we show that widely adopted $L_{1}$ penalization-based techniques for network pruning do not yield expected results. Finally, we explore the Edge of Stability phenomenon, indicating its inapplicability even to Lipschitz continuous convex differentiable functions, leaving its relevance to non-convex non-differentiable neural networks inconclusive. Our analysis exposes misguided interpretations of NGDMs in widely referenced papers and texts due to an overreliance on strong smoothness assumptions, emphasizing the necessity for a nuanced understanding of foundational assumptions in the analysis of these systems.
Machine Learning, ICML
Gradient Descent (GD) and its variants (Duchi et al., 2011; Kingma & Ba, 2014; Lydia & Francis, 2019; McMahan & Streeter, 2010; Shi & Li, 2021; Tieleman et al., 2012; Zhang, 2018) have played a pivotal role in advancing image and language processing over the last two decades. These algorithms have found widespread use in neural network training because the are easy to implement, highly scalable, and iteratively approach stationary points, where the loss function’s gradient is zero (Bertsekas, 1997).
Despite being initially designed for differentiable loss functions, GDs are widely used to minimize the loss functions of non-differentiable neural networks. The loss functions of these non-differentiable neural networks takes a special form – they are differentiable almost everywhere. Consequently, when gradient descents are applied to these non-differentiable neural networks (which we refer to as non-differentiable gradient methods or NGDMs), they seldom encounter non-differentiable points during the training process. This trait, combined with recent findings that NGDMs converge to stationary points under mild regularity conditions (Bolte & Pauwels, 2021; Davis et al., 2020), has led standard references in the field to assume that non-differentiability in neural networks is inconsequential (see Section 6.1.5 of (Stevens et al., 2020), for instance).
Moreover, this assumption has sparked renewed interest in the dynamics of gradient descents amid complex loss landscapes (Ahn et al., 2022b; Dauphin et al., 2014; Reddi et al., 2019), with fewer studies explicitly addressing non-differentiability (Bertoin et al., 2021; Bolte & Pauwels, 2021; Davis et al., 2020). Additionally, it has led theoretical papers to develop results for continuously differentiable functions and then claim universality by running simulations on non-differentiable neural networks (Experiment 6 in Ahn et al. (2022b), Section 3.1 in Ma et al. (2022), Section 6 in Zhang et al. (2022) to state a few).
In our paper, we challenge this assumption, demonstrating that it leads to inaccuracies in several highly referenced papers and texts regarding training dynamics in neural networks. We analyze three different aspects of the training process and make the following contributions:
  • Convergence Analysis of ReLU Networks: We show that the rate of convergence of NGDMs is slower than that of GDs. Our findings strongly suggest that the many claims about rate of convergence made under the assumption of $L-smoothness$ have little bearing on the properties of NGDMs.
  • Solution to the LASSO problem: We show that NGDMs do not produce sparse solutions for the LASSO problem even for the simplest case of the $L_{1}$ penalized linear model. Furthermore, we show that the NGDM solution to the LASSO problem can produce a counterintutive result – increasing the LASSO penalty can result in a solution with a larger $L_{1}$ norm.
  • The Edge of Stability: We demonstrate that the edge of stability conjecture does not hold, even for gradient descents on all convex differentiable neural networks. Furthermore, we provide a counter-example to Ma et al. (2022)’s claim that subquadratic behavior around the local minimum is responsible for the Edge of Stability phenomenon.
In a broader context, our analysis emphasizes the critical role the assumptions we make play on our understanding of these complex systems. For NGDMs in particular, our results suggest a need to treat GDs and NGDMs as distinct systems with different dynamics, challenging the common assumption of their similarity.
1 Preliminaries
In the three sections that follow, we show how disregarding the non-differentiability of the loss function severely limits our understanding of different aspects of training dynamics. Each of these sections is largely self-contained, complete with a literature review, problem setup, and analysis. Throughout this paper, experiments involving deep networks consistently employ convolutional neural network (CNN) architectures akin to VGG16 (Simonyan & Zisserman, 2014). This choice is motivated by the architecture’s widespread usage in the papers referenced throughout our discussions. The present section addresses background and terminology that is pertinent to all the subsequent sections.
NDGM v/s GD:
The crux of our analyses lies in the structural differences between GDs and NDGMs. For a continuously differentiable loss function, $f(\beta)$, the well-studied GD sequence (Bertsekas, 1997; Boyd et al., 2004) is described by the recursion
$\beta_{t+1}=\beta_{t}-\alpha\nabla f(\beta_{t}),$ (1)
where $\alpha$ is the learning rate, and $\beta_{0}$ is a randomly initialized starting point.
For a continuous non-differentiable loss function, $g(\beta)$, the NDGM recursion is given by
$\tilde{\beta}_{t+1}=\tilde{\beta}_{t}-\alpha\widetilde{\nabla}g(\tilde{\beta}_{t}),$ (2)
where ~$\widetilde{\nabla}g(\tilde{\beta}_{t})=\nabla g(\tilde{\beta}_{t})$ if the gradient exists at $\tilde{\beta}_{t}$, else it is a heuristic measure  [^1] https://rb.gy/g74av. The deep learning literature widely assumes that the dynamics described by the recursions in (1) and (2) are nearly the same.
Convex non-differentiable loss functions:
In this paper, we show the problems with this assumption using a variety of convex loss functions. The majority of our analysis is performed on the penalized single layer neural network regression problem with a non-positive response vector, which Kumar (2023) has shown to be convex. The loss function for this problem is given by
$\begin{split}L_{1}(\beta;{\bf{Z}},{\bf{y}},\lambda_{1},\lambda_{2})&=||{\bf{y}}-\max(0,{\bf{Z}}\beta)||^{2}\\ &=+\lambda_{1}||\beta||_{1}+\lambda_{2}||\beta||_{2}^{2},\end{split}$ (3)
where $\beta$ is a $P\times 1$ parameter vector, ${\bf{Z}}$ is an $N\times P$ data matrix whose $i^{th}$ column is given by $z_{i}$, ${\bf{y}}$ is a $N\times 1$ non-positive response vector, $\lambda_{1}\geq 0$ is the LASSO penalty, and $\lambda_{2}\geq 0$ is the ridge penalty. Throughout this paper, we denote the learning rate by $\alpha$, and the $i^{th}$ entry in the parameter vector after $k$ iterations is $\beta_{k}[i]$, with $10\leq i\leq P-$ 1.
The subgradient method:
Our analysis heavily relies on the observation that for a convex non-differentiable loss functions, the NDGM recursion described in (2) effectively implements the subgradient method (Bertsekas, 1997; Shor, 2012). The theoretical properties of this method are well studied in the optimization literature, with a concise overview of key findings available in the lecture notes of Boyd et al. (2003) and Tibshirani (2015). The subgradient method and GD recursions have different properties, and these differences play a key role in our subsequent discussions. We reference the relevant portions from the lecture notes as they arise.
Constant, diminishing, and reducing step sizes:
Our formulations in equations (1) and (2) employ a constant learning rate. This choice bears some explaining, for which we first recall some definitions from the optimization literature (see Chapter 1 of (Bertsekas, 1997) for details). In the GD literature, the training regimen described in equation (1) is called a constant step-size regime. In this regime, the GD is initialized using some randomized approach, and iterated with a constant learning rate until the loss stops decreasing. Although constant step-size GDs approach the stationary point, they rarely converge to it except in special cases.
To ensure provable convergence to the stationary point, theoretical studies often adopt the diminishing step-size regime. This regime is characterized by the iteration
$\beta_{t+1}=\beta_{t}-\alpha_{t}\nabla f(\beta_{t}),$ (4)
where $\lim_{t\rightarrow\infty}\alpha_{t}=0$, and $\sum_{t}\alpha_{t}\rightarrow\infty$. Convergence results in this regime are derived under ideal conditions, and provide guarantees that the recursion will reach the stationary point at least in the theoretical limit.
In practice, diminishing step-size GDs are implemented with a stopping criteria which determines when the solution produced by the GD is good enough to end training. The combination of the diminishing step-size regime and stopping criteria commonly used in the neural network literature is what we term the “reducing” step-size regime. In this regime, training unfolds as follows: commencing with a weight of $\beta_{0}$, GD is run for some $N_{0}$ iterations with a learning rate of $\alpha_{0}$, followed by $N_{1}$ iterations with a reduced learning rate of $\alpha_{1}<\alpha_{0}$, and so on, until the engineer is satisfied with the solution. Essentially, the reducing step-size regime involves chaining together a finite number (say $T$) of constant step-size GD runs, with the $(i+1)^{th}$ run having an initialization of $\beta_{u}$ and a learning rate of $\alpha_{i}$, where $u=\sum_{j=0}^{i-1}N_{j}$. In the reducing step size regime, the theoretical guarantees described in the diminishing step-size regime may not be realized. Nevertheless, the limiting dynamics of the sequence in this regime are exactly those of a constant step-size GD, with learning rates and weight initializations chosen corresponding to the values at the beginning of the $T^{th}$ run. Accordingly, in this paper, we exclusively study the constant step-size problem, assuming that the weights and learning rates have been appropriately chosen.
2 NDGM and convergence analysis
Recent theoretical work by Davis et al. (2020) has shown that in the diminishing step-size regime, NDGMs converge to a local optimal solution in the theoretical limit under fairly general conditions. Unfortunately, these regularity conditions are rarely satisfied in practice. To illustrate this point, consider the Assumption A.3 in Davis et al. (2020), which requires that in the diminishing step size regime, the sum of the learning rates should diverge. As was previously stated, this is a standard assumption even in the thoeretical convergence analysis of GDs. However, the prevelant step-size reduction we observed in most training scripts involves decaying the learning rate by a constant multiplicative factor at predefined intervals (StepLR in PyTorch [^2] https://rb.gy/pub4s). With this decay scheme, the learning rates form a convergent geometric series, thus violating the assumption.
For the differentiable case, we know that even when the learning rates form a convergent geometric series, the GD sequence will come within some vicinity of the optimal solution. Therefore, the question arises if the same holds true for NDGM sequences. More precisely, can we state that the dynamics of NDGM and GD sequences are identical?
NDGM and GD dynamics are not identical:
The dynamics of the two sequences are not identical, and we demonstrate this by arriving at a contradiction. Suppose the sequence $\{\beta_{k}\}_{k=0}^{\infty}$ obtained by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})=L_{1}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1},0,0.01)$ has the same properties as a GD sequence. Then, the entire NDGM sequence can be characterized using the ”Capture Theorem” as described below
Proposition 2.1.
Let $x_{k+1}=x_{k}-\alpha\nabla f(x_{k})$ be a sequence generated by running GD on a continuously differentiable function, $f(x)$ such that $f(x_{k+1})\leq f(x_{k})$. If $x^{*}$ is an isolated local minimum of $f(x)$ in a unit ball around $x^{*}$, and $||x_{0}-x^{*}||<1$, then $||x_{k}-x^{*}||<1$ for all $k>0$.
Proof.
Result follows by using $\bar{\epsilon}=1$ in Proposition 1.2.5 of (Bertsekas, 1997). ∎
Therefore, if $||\beta_{0}||_{\infty}<1$, and $L_{2}(\beta_{k+1};{\bf{Z}}_{1},{\bf{q}}_{1})\leq L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$ for all $k\geq 0$, then we expect $||\beta_{k}||_{\infty}<1$ for all $k\geq 0$. To validate this claim, we first generate a $20\times 500$ matrix ${\bf{Z}}_{1}$ with entries uniformly sampled on $[-1,1]$. Next, we generate a $20\times 1$ vector ${\bf{q}_{1}}$ with entries uniformly sampled on $[-1,0]$. Finally, we generate the sequence $\{\beta_{t}\}_{t=0}^{\infty}$, by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})$ with the entries in $\beta_{0}$ uniformly sampled on $[-1,1]$. As seen in Figure 1, although the loss function is a monotone decreasing function of the iteration number, the infinity norm of the NDGM sequence eventually takes values greater than 1, hence establishing the contradiction. The non-capturing described in Figure 1 is to be expected; as stated in Section 1 of Boyd et al. (2003), ”Unlike the ordinary gradient method, the subgradient method is not a descent method; the function value can (and often does) increase”.
Figure 1: Plot of the loss function ($L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$), and infinity norm ($||\beta_{k}||_{\infty}$) of the parameter vector as a function of the iteration number ($k$). Although the loss function is monotone decreasing, the infinity norm of the weights does not ”capture”, hence demonstrating that the dynamics of NDGM and GD are not identical
Convergence analysis and L-smoothness:
Given that the dynamics of GDs and NDGMs differ, the question arises what assumptions are appropriate for the study of convergence analysis in neural network training. Majority of the analyses in the literature (Défossez et al. (2020); Reddi et al. (2019); You et al. (2019); Zhang et al. (2022) to state a few) assume that the loss function, $f(\beta)$ is continuously differentiable, and is $L-smooth$, i.e., there exists a constant $L$, such that
$||\nabla f(x)-\nabla f(y)||\leq L||x-y||$ (5)
for all $x$, $y$ in the domain of $f$. However, this assumption breaks down for non-differentiable functions, where the gradient does not even exist at various points in the domain. This leads us to a pertinent inquiry: Can we salvage this assumption by confining its validity to a local context, possibly within a specific interval centered around the local optimal solution?
Unfortunately, even this cannot be done; as stated in (Laurent & Brecht, 2018), “local minima of ReLU networks are generically nondifferentiable. They cannot be waved away as a technicality, so any study of the loss surface of such network must invoke nonsmooth analysis”. Moreover, assuming that the loss function is $L-smooth$ grossly over-estimates the speed of convergence of NDGMs. To illustrate this, consider that in order to achieve an error level of $\epsilon$, GD applied to an $L-smooth$ convex loss function, $f$, would require $O(1/\epsilon)$ steps for convergence. In contrast, NDGM applied to a convex non-differentiable loss function, $g$, satisfying the Lipschitz condition
$||g(x)-g(y)||\leq L||x-y||$ (6)
would take $O(1/\epsilon^{2})$ steps (see Section 7.1.3 of Tibshirani (2015) for the result, and a visual illustration of how much slower $O(1/\epsilon^{2})$ is than $O(1/\epsilon)$). Thus, it is likely that the vast literature on the convergence analysis of GDs for loss functions satisfying (5) likely have little bearing on the dynamics of convergence in modern non-smooth neural networks. More generally, we expect the speed of convergence for smooth neural networks to be faster than that of non-smooth neural networks.
Experiments with NDGM convergence:
To validate our intuition in a more general context, we conduct experiments on CNNs of various depths (see the appendix for architecture details). For each network, we carry out two identical simulations, with the only difference being the activation function used. In the first simulation, we employ the non-smooth ReLU activation function, and in the second, we use the similarly shaped, but smooth and asymptotally equivalent, GELU activation function (Hendrycks & Gimpel, 2016).
As seen in Figure 2, the difference in convergence speeds between smooth and non-smooth networks widens as the depth of the neural network. This observation aligns with our expectations, given that increased depth introduces greater non-smoothness into the system. Hendrycks & Gimpel (2016) substantiate a similar trend of accelerated convergence in GELU networks compared to ReLU networks across diverse architectures and datasets (see Figures 2 through 7 in their paper).
Figure 2: Training curves comparing smooth and non-smooth CNNs (from left to right) of depth 11, 13, and 19 respectively. As the network depth increases, the non-smoothness in the system also increases. Consequently, the difference in convergence speeds between smooth and non-smooth networks widens with the increasing non-smoothness in the system.
Do results derived in the L-Smooth setting describe the behavior of all differentiable neural networks?:
In a subsequent section on the Edge of Stability, we show that insights based on the assumption of $L-smoothness$ are not universally applicable even across all Lipschitz continuous, convex, differentiable neural networks. We point out that $L-smoothness$ is one of the strongest assumptions which can be applied to a function after convexity. This assumption guarantees the existence of the first and second derivatives, and the boundedness of the second derivative—a collection of criteria that are unmet by the majority of the loss functions encountered in the neural network literature.
The way forward:
In this section, we show that NDGMs are slower to converge than GDs. Fortunately, most recent explorations into image and language models primarily leverage the transformer architecture (Vaswani et al., 2017) which employs the differentiable GELU activation function (Hendrycks & Gimpel, 2016). As a consequence, future iterations of neural networks may predominantly employ pure GDs. Therefore, it is crucial for the research community to establish a consensus on a set of assumptions that accurately represent these complex systems before meaningful progress can be achieved.
3 NDGM and the LASSO penalty
Zero weights play a critical role in model compression (Blalock et al., 2020; Han et al., 2015a, b; Li et al., 2016), and it is desirable to force weights down to zero whenever we can do so without sacrificing too much accuracy. A common approach to induce sparsity is to penalize the $L_{1}$ norm of the model weights, with the penalizing factor commonly referred to as the LASSO penalty. In the statistics literature, loss functions involving penalized $L_{1}$ norms are typically minimized using specialized algorithms designed for different use cases (Efron et al., 2004; Friedman et al., 2008; Mazumder & Hastie, 2012; Tibshirani, 1996; Tibshirani et al., 2005; Zou & Hastie, 2005). However when training neural networks, the penalized $L_{1}$ norm is used with arbitrary non-convex functions, with the understanding that running NDGM on the problem results in a near sparse optimal solution (Bengio, 2012; Goodfellow et al., 2016). As stated in Scardapane et al. (2017), ”Even if (the LASSO) has a non differentiable point in 0, in practice this rarely causes problems to standard first- order optimizers. In fact, it is common to simultaneously impose both weight-level sparsity with the $L_{1}$ norm, and weight minimization using the $L_{2}$ norm, resulting in the so-called elastic net penalization”. This expectation is not being met in reality, and there is strong evidence of users struggling with the interpretation of their LASSO results in the PyTorch forums  [^3] https://rb.gy/kh0qr  [^4] https://rb.gy/2vsry  [^5] https://rb.gy/6dfli. Given the wide variety of critical applications in which the LASSO finds use (Ghosh et al. (2021); Li et al. (2011); Li & Sillanpää (2012); Ogutu et al. (2012) to state a few), it is critical to understand the reliability of the NDGM solution for the LASSO problem.
NDGM solutions for the LASSO problem are not reliable:
We show that NDGM produces unreliable solutions even for the simple case of the LASSO penalized linear model. In particular, we show that increasing the LASSO penalty can result in a NDGM solution with a larger $L_{1}$ norm, completely defeating the purpose of the penalty. We first analytically demonstrate the unreliability for the simplest LASSO problem,
$L_{3}(\beta)=L_{1}(\beta;0,0,\lambda_{1},0)=\lambda_{1}||\beta||_{1},$ (7)
with $\lambda_{1}>0$, and then demonstrate the same issues for the general LASSO problem through simulations. We begin by showing that for any non-zero learning rate, $\alpha$, the NDGM sequence for (7) will be non-sparse with probability 1.
Proposition 3.1.
Let $\{\beta_{t}\}_{t=0}^{\infty}$ denote the sequence generated by running NDGM on (7), with a constant learning rate, $\alpha$, and the entries in $\beta_{0}$ being uniformly sampled from $[-1,1]$. Then $\beta_{k}$ will have all non-zero entries with probability 1 for all $k>0$.
Proof.
The NDGM iteration for (7) is described by the recursion
$\beta_{i}[k]=\begin{cases}\beta_{i-1}[k]-\alpha\lambda_{1}&\text{ if $\beta_{i-1}[k]>0$}\\ \beta_{i-1}[k]+\alpha\lambda_{1}&\text{ if $\beta_{i-1}[k]<0$}\\ 0&\text{ if $\beta_{i-1}[k]=0$}\end{cases}.$ (8)
From (8), it is clear that if the $k^{th}$ entry in the vector becomes zero during any iteration, then it stays zero for all subsequent iterations. Accordingly, suppose the $k^{th}$ entry in the NDGM sequence vector becomes 0 for the first time after $N$ iterations. Then we have
$\beta_{N}[k]=\beta_{0}[k]+(N-2m)\alpha\lambda_{1}=0,$ (9)
where $m$ is the number of times the $k^{th}$ entry of the parameter exceeds 0 in the first $N$ iterations. (9) can only hold if $\beta_{0}[k]$, is an integer multiple of $\alpha\lambda_{1}$. Since the feasible values are countable, and the set of initializations is uncountable, the probability of the occurrence has measure 0. ∎
In the next proposition, we show that the sequence does not converge, but oscillates between two fixed points.
Proposition 3.2.
The sequence $\{\beta_{t}\}_{t=0}^{\infty}$ described in the previous proposition does not converge. Furthermore, there is an integer $N_{0}$ such that for every $N\geq N_{0}$, the sequence oscillates between two vectors : $\gamma_{1}$ and $\gamma_{2}$, where $||\gamma_{1}||_{\infty}<\alpha\lambda_{1}$, and $||\gamma_{2}||_{\infty}<\alpha\lambda_{1}$.
Proof.
We prove the results assuming $\beta_{0}[k]>0$; the proof is similar when $\beta_{0}[k]<0$. From Proposition 3.1, we know that for any $N>0$, $\beta_{N}[k]\neq 0$ with probability 1. Therefore, (8) implies that starting from any $\beta_{0}[k]>0$, the sequence will decrease monotonically till it reaches a value between 00 and $\alpha\lambda_{1}$. Accordingly, let $\beta_{n_{k}}[k]=\gamma_{k}$ for some $n_{k}\geq 0$, with $\gamma_{k}\in(0,\alpha\lambda_{1})$. Since $\gamma_{k}-\alpha\lambda_{1}<0$, completing the recursion in (8) gives
$\beta_{n_{k}+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even}.\end{cases}$ (10)
The result follows by choosing $N_{0}=\text{max}(n_{0},n_{1}\ldots n_{P-1})$. ∎
Equation (10) suggests the surprising result that with the same learning rate and initialization, a larger value of $\lambda_{1}$ can result in a NDGM solution with a larger $L_{1}$ norm, which completely defeats the purpose of the LASSO penalty. To better understand the issue, consider the 2D case, with the weights initialized as $\beta_{0}=[0.5053,0.5053]$. With $\lambda_{1}=1$ and $\alpha=0.01$, Proposition 3.2 states that the NDGM sequence will oscillate between $[0.0053,0.0053]$ and $[-0.0047,-0.0047]$, but with $\lambda_{1}=100$, and the same learning rate and starting point, the solution will oscillate between $[0.5053,0.5053]$ and $[-0.4947,-0.4947]$. Thus, at the point when NDGM appears to converge, the solution with $\lambda_{1}=1$ will have an $L_{1}$ norm of at most 0.0106, but solution with $\lambda_{1}=100$ case will have an $L_{1}$ norm of atleast 0.9894. Furthermore, note that reducing the step-size does not resolve this paradox. If, in both cases, we decrease the learning rate to $\alpha=0.001$ once the algorithm seems to have converged, then with $\lambda_{1}=1$, the NDGD sequence will oscillate between $[0.0003,0.0003]$ and $[-0.0007,-0.0007]$. However, with $\lambda_{1}=100$, the NDGM sequence will oscillate between $[0.0053,0.0053]$ and $[-0.0947,-0.0947]$, and the paradox persists.
To demonstrate the same issue in the broader setting, consider the loss function for the general LASSO problem given by
$L_{4}(\beta;{\bf{W}},{\bf{y}},\lambda_{1})=\frac{1}{N}||{\bf{y}}-{\bf{W}}\beta||_{2}^{2}+\lambda_{1}||\beta||_{1},$ (11)
where ${\bf{W}}$ is an arbitrary $20\times 500$ dimensional data matrix, and ${\bf{y}}$ is an arbitrary $20\times 1$ response, each of whose entries are sampled from the uniform distribution on $[-1,1]$. We run NDGM on (11) twice with the same learning rate and initialization, once with $\lambda_{1}=0.01$, and once with $\lambda_{1}=10$. At the end of the two runs, the optimal solution with $\lambda_{1}=0.1$ has an $L_{1}$ norm of 1.62, and the optimal solution with $\lambda_{1}=10$ has an $L_{1}$ norm of 25.4 i.e., a 1000 fold increase in the value of $\lambda_{1}$ results in a more than 15 fold increase in the $L_{1}$ norm of the optimal solution!
Connections to the subgradient method:
The counter-intuitive behavior seen in the example above can be explained using the subgradient method. Since the loss function described in equation (7) satisfies the Lipschitz condition (see equation (6)) with $L=\lambda_{1}$, we have from Boyd et al. (2003) (Section 2 on the Constant Step Size) that
$\lim_{k\rightarrow\infty}||\beta_{k}||_{1}<\alpha\lambda_{1}.$ (12)
Therefore, for the same value of $\alpha$, the larger value of $\lambda_{1}$ will result in a larger error in estimation, which is precisely what we are seeing above.
Diminishing v/s reducing step sizes:
This illustration also highlights the contrast between the diminishing and reducing step size regimes outlined in the preliminaries. Under the diminishing step-size regime, $\alpha\rightarrow 0$ in the theoretical limit, and therefore, Proposition 3.1 guarantees that the optimal solution for (7) will be 0 for all $\lambda_{1}>0$. On the other hand, in the reducing step-size regime, $\alpha>0$ at the end of training, and therefore, the error in estimation becomes a critical determinant of the optimal value.
Experiments with deep networks:
In their seminal work on network pruning, Han et al. (2015b) propose the utilization of $L_{1}$ regularization to encourage weights to approach 0, facilitating the pruning process. This approach is widely adopted in the literature (Wen et al., 2016; Xiao et al., 2019; Yang et al., 2019), and is explicitly endorsed in NVIDIA’s guides for training standard object detection models such as RetinaNet and YOLOv4666See the recommendation of the regularizer in the Training config section of https://rb.gy/20y13m and https://rb.gy/i1sb6s. To challenge the general applicability of this suggestion, we fine-tune two pretrained VGG16 models using the $L_{1}$-penalized cross entropy loss with a learning rate of $\alpha=0.01$, with one model using a regularization parameter of $\lambda_{1}=0.001$ and the other with $\lambda_{1}=0.1$. At the end of fine-tuning, the $L_{1}$ norm of the weights in the model fine-tuned with $\lambda_{1}=0.1$ is more than 30 times larger (69178.44) than that of the model fine-tuned with $\lambda_{1}=0.001$ (2113.83). Furthermore, the latter model has 99.7% of its weights less than $10^{-5}$, whereas the former model has only 0.01% of its weights less than $10^{-5}$, further underscoring the discrepency. We conducted experiments with different values of $\alpha$, and the widely used Adam optimizer (Kingma & Ba, 2014), and in each case, we observed the same paradox. Our code is provided in the appendices.
NDGM and the way forward with sparsity:
While NDGMs do not inherently generate sparse solutions, iterative methods have been developed to produce sparse solutions for the $L_{1}$ regularized linear problem (see Table 1 in Bottou (2010)), and $L_{1}$ regularized neural networks in general (Ma et al., 2019; Siegel et al., 2020). The realization that NDGMs lack the capability to produce sparse solutions predates the widespread adoption of deep learning. As mentioned in (Xiao, 2009), “… stochastic gradient descent (SGD), has limited capability of exploiting problem structure in solving regularized learning problems”. Despite this, a prevalent misconception persists in the deep learning literature, erroneously suggesting that such algorithms yield sparse solutions, and that the solutions generated by deep learning frameworks like PyTorch (Paszke et al., 2017) ought to be comparable to those generated using traditional machines learning frameworks like glmnet (Friedman et al., 2021) or scikit-learn (Kramer & Kramer, 2016). We point out that there exits a fundamental disparity between traditional machine learning frameworks like glmnet and deep learning frameworks like PyTorch. glmnet focuses on a narrow set of optimization problems, concealing implementation intricacies while guaranteeing optimality. Conversely, PyTorch provides greater flexibility in constructing models but lacks similar guarantees, placing the responsibility for solution quality on users. Attaining a solution at a non-differentiable point poses a challenging problem, underscoring the significance of exploring related studies and tools rather than opting for a simplistic implementation, particularly when precision is paramount.
Connections to the Edge of Stability:
Proposition 3.2 shows that the NDGM sequence for the non-differentiable, Lipschitz continuous, convex loss function described in (7) will not diverge to $\infty$ for any finite value of $\alpha$. This is a special case of unstable convergence, a topic we cover in greater detail in the next section.
4 NDGM and the Edge of Stability
In recent years, there has been a growing interest on the topic of “unstable convergence”. The literature on the subject (Ahn et al., 2022a, b; Arora et al., 2022; Chen & Bruna, 2022; Cohen et al., 2021; Li et al., 2022) is motivated by the finding that unlike convex quadratic forms, gradient descent on neural networks do not diverge to $\infty$ even when the learning rate is greater than $\alpha^{*}=2/\eta$, where $\eta$ is the dominant eigenvalue of the loss function’s Hessian. Instead, the loss function has been empirically demonstrated to converge “unstably”, reducing non-montonically on the long run. $\alpha=\alpha^{*}$ is referred to in the literature as the ”Edge of Stability”, since it separates regions of ”stable convergence” ($\alpha<\alpha^{*}$) from regions of “unstable convergence” [^7] We define unstable convergence to mean “non-divergence to $\infty$”. Our definition is adapted from Definition 1.1 of Arora et al. (2022). However, we deviate from their definition due to its dependence on the existence of the Hessian at all points between the current and next iterate, a condition that need not be satisfied in non-smooth neural networks. ($\alpha>\alpha^{*}$).
The derivation of the Edge of Stability condition assumes that the loss function, $f$, is $L-smooth$ (see equation (5)). Since the result has been demonstrated in most convex optimization settings (Jacot et al., 2018; Li & Liang, 2018), it is conjectured that the result ought to hold even in the case of non-convex, non-differentiable neural networks. This conjecture is incorrect even for Lipschitz continuous, differentiable, convex functions satisfying (6) as shown in the proposition below.
Proposition 4.1.
All convex non-smooth loss functions having bounded subgradients, or satisfying equation (6) will show unstable convergence.
Proof.
Note that any loss function satisfying (6) has subgradients bounded above by $L$. Additionally, as outlined in Section 1 of Boyd et al. (2003), the deductions derived from the subgradient method hold equivalently for differentiable functions. This equivalence arises from the fact that the only admissible value for the subgradient of a differentiable convex function at any given point is, indeed, its gradient. Accordingly, using the convergence properties of the subgradient methods with constant step sizes from Section 2 of Boyd et al. (2003), we have that $\lim_{k\rightarrow\infty}f(x_{k})-f^{*}\leq\alpha L^{2}<\infty$, hence the result. ∎
Figure 3: Training curves for $\alpha=0.1$, $\alpha=1$, and $\alpha=10$ with the Lipschitz continuous Huber loss function. As the learning rate increases, the loss starts oscillating with larger frequencies, but never diverges to $\infty$. This is to be expected from Proposition 4.1, since the Huber loss function is convex, non-smooth and has bounded subgradients.
In the previous section, we analytically derive the unstable convergence for the LASSO problem described in (7). An example of a differentiable non-smooth convex function satisfying Proposition 4.1 is the Huber loss function, which finds use in many state of the art object detection models (Girshick, 2015; Liu et al., 2016; Ren et al., 2015). The Huber loss function is non-smooth because it is once but not twice differentiable. For a regression problem involving an arbitrary $50\times 1$ response vector, $\mathbf{y}=[y_{1},y_{2}\ldots y_{N}]^{T}$, and an arbitrary $50\times 200$ data matrix, $\mathbf{Z}$, the Huber loss function is defined as
$L_{7}(\beta)$ $=\frac{1}{N}\sum l(i),\quad\text{where}$ (13)
$l(i)$ =$=\begin{cases}\frac{1}{2}(y_{i}-z_{i}^{T}\beta)^{2}&\text{if }\left|(y_{i}-z_{i}^{T}\beta)\right|<1\\ \left(|y_{i}-z_{i}^{T}\beta|-\frac{1}{2}\right)&\text{otherwise}\end{cases}$
where $z_{i}$ is the $i^{th}$ column of $\mathbf{Z}$. Figure 3 shows that the NDGM sequence for (13) does not diverge toward $\infty$, even with a high learning rate, such as $\alpha=10$, which aligns with our expectations. Furthermore, the figure reveals an intriguing loss pattern when $\alpha=1$: the loss initially increases before the oscillations become noticeable. This peculiar behavior has been documented in the literature (Ahn et al., 2022b; Arora et al., 2022), although a conclusive explanation for this phenomenon remains elusive at this time.
Proposition 4.1 suggest that it is unclear and the examples that follow show that the common approach of extrapolating results developed for $L-smooth$ loss functions to the non-smooth or non-differentiable case is problematic. In the next subsection, we demonstrate the problems with this approach by providing an example of an extrapolation which breaks down in the non-smooth case.
Subquadratic losses and unstable convergence
In section 3.1 of Ma et al. (2022), the authors employ second-order finite differences to investigate the curvature of VGG16 and ResNet in the vicinity of the local minimum. They conclude that subquadratic growth of the loss function around this minimum point explains the edge of stability phenomenon. However, their analysis is flawed because VGG16 and ResNet, as ReLU networks, have local minima that are generically non-differentiable (Laurent & Brecht, 2018). As a result, the points of interest lie where their approximation breaks down.
To challenge the broad applicability of their statement, consider the Huber loss function depicted in Equation (13). This function is overparameterized and therefore possesses a global minimum of 0. Consequently, within the vicinity of this global minimum, the loss function exhibits quadratic growth. In such instances, Ma et al. (2022)’s result states that we will not see unstable convergence. However, contrary to their suggestions, Proposition 4 guarantees unstable convergence under these circumstances, as evidenced in Figure 3.
The discrepancy arises because the Huber loss function lacks the twice differentiability required by Definition 4.1 of Ma et al. (2022)’s paper. In contrast, all $L-smooth$ functions will exhibit subquadratic growth near their local minimum during unstable convergence, as noted by (Ma et al., 2022). This occurs because $L-smooth$ functions are twice differentiable by definition. This example highlights the challenges of extending results developed under strong assumptions like $L-smoothness$ to the general case.
The way forward:
We demonstrate that our understanding of the Edge of Stability will be incomplete until we more precisely quantify conditions under which we expect it to occur. Furthermore, considering ReLU networks’ non-differentiable minima, without further research into NDGM behavior around non-differentiable points our grasp of training dynamics lacks completeness. Consequently, we propose employing the convex non-differentiable single-layer neural network presented in (Kumar, 2023) as the toy problem for theory development. Studying this neural network offers several advantages: 1) its convex nature renders it more amenable to analysis compared to non-convex functions, 2) we possess analytical knowledge of the global optimum, and 3) the global optimum exists at a point of non-differentiability—the setup which is particularly relevant to ReLU networks.
5 Conclusion
In this paper, we highlight how non-differentiability alters training dynamics across three domains: convergence analysis, the LASSO problem, and the Edge of Stability. In each case, we show that problems arise because of incorrect assumptions made about the contour of the loss surface. Our results suggest the critical need to reach consensus on a set of assumptions which can be made about the loss surface before progress can be made.
Appendix A Appendix A - Architecture and code for Figure 2
The code to construct our plots in Figure 2, was taken from https://github.com/kuangliu/pytorch-cifar/blob/master/models/vgg.py, with three minor modifications. First, we removed the batch normalization as part of each layer. We did this because, while it is known that batch normalization alters the training dynamics (Ioffe & Szegedy, 2015), and our work is interested in the effect of network smoothness on convergence speed in the absence of confounders. Second, we replaced max pool with average pool to ensure differentiability of the layer, and third, we used a simpler layer configuration than one provided in the link. The configuration we used is provided below:
cfg = {
    ’VGG11’: [16, ’M’, 32, ’M’, 32, 32, ’M’, 64, 64, ’M’, 64, 64, ’M’],
    ’VGG13’: [16, 16, ’M’, 32, 32, ’M’, 32, 32, ’M’, 64, 64, ’M’, 64, 64, ’M’],
    ’VGG19’: [16, 16, ’M’, 32, 32, ’M’,
              32, 32, 32, 32, ’M’,
              64, 64, 64, 64, ’M’,
              64, 64, 64, 64, ’M’],
Appendix B Appendix B - Code for paradoxes relating to the LASSO
The training code was generated using ChatGPT and verified for correctness. The code below utilizes SGD with a learning rate of 0.01, though can trivially be changed to any other.
import torch
import torch.nn as nn
import torch.optim as optim
from torchvision import datasets, transforms, models
# Set random seed for reproducibility
seed = 42
torch.manual_seed(seed)
torch.backends.cudnn.deterministic = True
torch.backends.cudnn.benchmark = False
import numpy as np
np.random.seed(seed)
# Set device to GPU if available
device = torch.device("cuda" if torch.cuda.is_available() else "cpu")
# Define transform and download CIFAR-10 dataset
transform = transforms.Compose([
    transforms.ToTensor(),
    transforms.Normalize((0.5, 0.5, 0.5), (0.5, 0.5, 0.5)),
train_dataset = datasets.CIFAR10(root=’./data’, train=True, download=True, transform=transform)
train_loader = torch.utils.data.DataLoader(train_dataset, batch_size=512, shuffle=True, num_workers=4)
# Load pretrained VGG16 model
model = models.vgg16(pretrained=True)
model.to(device)
# Define L1 regularization strength
l1_strength = 0.1
# Define the loss function (CrossEntropyLoss + L1 penalty)
criterion = nn.CrossEntropyLoss()
optimizer = optim.SGD(model.parameters(), lr=0.01)
# Training loop
epochs = 150
for epoch in range(epochs):  # Adjust the number of epochs as needed
    for inputs, labels in train_loader:
        inputs, labels = inputs.to(device), labels.to(device)
        # Forward pass
        outputs = model(inputs)
        loss_ce = criterion(outputs, labels)
        # L1 penalty
        l1_penalty = 0
        for param in model.parameters():
            l1_penalty += torch.abs(param).sum()
        # Add L1 penalty to the loss
        loss = loss_ce + l1_strength * l1_penalty
        # Backward and optimize
        optimizer.zero_grad()
        loss.backward()
        optimizer.step()
    print(f’Epoch [{epoch+1}/100], L1 Penalty: {l1_penalty.item():.4f}’)
print("Training complete!")
total_params = 0
num_small_params = 0
threshold = 1e-5
for name, param in model.named_parameters():
    total_params += param.numel()
    if ’weight’ in name or ’bias’ in name:
        num_small_params += torch.sum(torch.abs(param) < threshold).item()
fraction_small_params = num_small_params / total_params
print(f’Fraction of parameters with values less than 10^-5: {fraction_small_params:.5f}’)

### v3

@@
 Consequently, we show that widely adopted $L_{1}$ penalization-based techniques for network pruning do not yield expected results.
+Additionally, we dispel the common belief that optimization algorithms like Adam and RMSProp perform similarly in non-differentiable contexts.
 Finally, we explore the Edge of Stability phenomenon, indicating its inapplicability even to Lipschitz continuous convex differentiable functions, leaving its relevance to non-convex non-differentiable neural networks inconclusive.
@@
 This trait, combined with recent findings that NGDMs converge to stationary points under mild regularity conditions (Bolte & Pauwels, 2021; Davis et al., 2020), has led standard references in the field to assume that non-differentiability in neural networks is inconsequential (see Section 6.1.5 of (Stevens et al., 2020), for instance).
-Moreover, this assumption has sparked renewed interest in the dynamics of gradient descents amid complex loss landscapes (Ahn et al., 2022b; Dauphin et al., 2014; Reddi et al., 2019), with fewer studies explicitly addressing non-differentiability (Bertoin et al., 2021; Bolte & Pauwels, 2021; Davis et al., 2020).
-Additionally, it has led theoretical papers to develop results for continuously differentiable functions and then claim universality by running simulations on non-differentiable neural networks (Experiment 6 in Ahn et al.
-(2022b), Section 3.1 in Ma et al.
-(2022), Section 6 in Zhang et al.
-(2022) to state a few).
+To emphasize this point, we note that three of the most widely cited surveys on optimization methods for neural networks (Bottou et al., 2018; Le et al., 2011; Ruder, 2016) do not mention non-differentiability, even though the ReLU, arguably the most common activation function in neural networks, is non-differentiable.
+Moreover, this assumption has sparked renewed interest in the dynamics of gradient descents amid complex loss landscapes ((Ahn et al., 2022b; Dauphin et al., 2014; Reddi et al., 2019) to state a few), with fewer studies explicitly addressing non-differentiability (Bertoin et al., 2021; Bolte & Pauwels, 2021; Davis et al., 2020).
+Additionally, it has led theoretical papers to develop results for continuously differentiable functions and then claim universality by running simulations on non-differentiable neural networks (Experiment 6 in (Ahn et al., 2022b), Section 3.1 in (Ma et al., 2022), Section 6 in (Zhang et al., 2022) to state a few).
 In our paper, we challenge this assumption, demonstrating that it leads to inaccuracies in several highly referenced papers and texts regarding training dynamics in neural networks.
@@
 Our findings strongly suggest that the many claims about rate of convergence made under the assumption of $L-smoothness$ have little bearing on the properties of NGDMs.
-  • Solution to the LASSO problem: We show that NGDMs do not produce sparse solutions for the LASSO problem even for the simplest case of the $L_{1}$ penalized linear model.
-Furthermore, we show that the NGDM solution to the LASSO problem can produce a counterintutive result – increasing the LASSO penalty can result in a solution with a larger $L_{1}$ norm.
+Furthermore, we dispel the widespread belief that different NDGM variants like RMSProp and Adam largely differ only in their rates of convergence, but converge to roughly the same eventual point in the limit.
+  • Solution to the LASSO problem: We show that NGDMs, including variants like RMSProp and NDGM with momentum, do not produce sparse solutions for the LASSO problem, even for the simplest case of the $L_{1}$ penalized linear model, and exhibit different behaviors, contrary to prevalent wisdom.
+Moreover, NGDMs can yield counterintuitive results, such as producing a solution with a larger $L_{1}$ norm with a larger LASSO penalty.
   • The Edge of Stability: We demonstrate that the edge of stability conjecture does not hold, even for gradient descents on all convex differentiable neural networks.
-Furthermore, we provide a counter-example to Ma et al.
-(2022)’s claim that subquadratic behavior around the local minimum is responsible for the Edge of Stability phenomenon.
+Furthermore, we provide a counter-example to the claim made in (Ma et al., 2022) that subquadratic behavior around the local minimum is responsible for the Edge of Stability phenomenon.
 In a broader context, our analysis emphasizes the critical role the assumptions we make play on our understanding of these complex systems.
@@
 This choice is motivated by the architecture’s widespread usage in the papers referenced throughout our discussions.
+Our experiments have relatively low computational costs, with the most expensive run completing within 2 hours on a single GPU.
 The present section addresses background and terminology that is pertinent to all the subsequent sections.
@@
 In this paper, we show the problems with this assumption using a variety of convex loss functions.
-The majority of our analysis is performed on the penalized single layer neural network regression problem with a non-positive response vector, which Kumar (2023) has shown to be convex.
+The majority of our analysis is performed on the penalized single layer neural network regression problem with a non-positive response vector, which (Kumar, 2023) has shown to be convex with with the all-zeros vector as its unique global minimizer.
 The loss function for this problem is given by
 $\begin{split}L_{1}(\beta;{\bf{Z}},{\bf{y}},\lambda_{1},\lambda_{2})&=||{\bf{y}}-\max(0,{\bf{Z}}\beta)||^{2}\\ &=+\lambda_{1}||\beta||_{1}+\lambda_{2}||\beta||_{2}^{2},\end{split}$ (3)
-where $\beta$ is a $P\times 1$ parameter vector, ${\bf{Z}}$ is an $N\times P$ data matrix whose $i^{th}$ column is given by $z_{i}$, ${\bf{y}}$ is a $N\times 1$ non-positive response vector, $\lambda_{1}\geq 0$ is the LASSO penalty, and $\lambda_{2}\geq 0$ is the ridge penalty.
-Throughout this paper, we denote the learning rate by $\alpha$, and the $i^{th}$ entry in the parameter vector after $k$ iterations is $\beta_{k}[i]$, with $10\leq i\leq P-$ 1.
+where $\beta$ is a $P\times 1$ parameter vector, ${\bf{Z}}$ is an $N\times P$ data matrix whose $i^{th}$ column is given by $z_{i}$, ${\bf{y}}$ is a $N\times 1$ non-positive response vector, $\lambda_{1}\geq 0$ is the LASSO penalty, $\lambda_{2}\geq 0$ is the ridge penalty, and $\lambda_{1}+\lambda_{2}>0$.
+Throughout this paper, we denote the learning rate by $\alpha$, and the $i^{th}$ entry in the parameter vector after $k$ iterations as $\beta_{k}[i]$, with $10\leq i\leq P-$ 1.
 The subgradient method:
 Our analysis heavily relies on the observation that for a convex non-differentiable loss functions, the NDGM recursion described in (2) effectively implements the subgradient method (Bertsekas, 1997; Shor, 2012).
-The theoretical properties of this method are well studied in the optimization literature, with a concise overview of key findings available in the lecture notes of Boyd et al.
-(2003) and Tibshirani (2015).
+The theoretical properties of this method are well studied in the optimization literature, with a concise overview of key findings available in the lecture notes of (Boyd et al., 2003) and (Tibshirani, 2015).
 The subgradient method and GD recursions have different properties, and these differences play a key role in our subsequent discussions.
@@
 In this regime, the GD is initialized using some randomized approach, and iterated with a constant learning rate until the loss stops decreasing.
-Although constant step-size GDs approach the stationary point, they rarely converge to it except in special cases.
-To ensure provable convergence to the stationary point, theoretical studies often adopt the diminishing step-size regime. This regime is characterized by the iteration
+Although constant step-size GDs approach the stationary point, they rarely converge to it unless the loss function possesses desirable properties like $L-smoothness$ (Proposition 1.2.3 in (Bertsekas, 1997)).
+To ensure provable convergence to the stationary point in more general settings, studies often adopt the diminishing step-size regime.
+This regime is characterized by the iteration
 $\beta_{t+1}=\beta_{t}-\alpha_{t}\nabla f(\beta_{t}),$ (4)
@@
 In this regime, training unfolds as follows: commencing with a weight of $\beta_{0}$, GD is run for some $N_{0}$ iterations with a learning rate of $\alpha_{0}$, followed by $N_{1}$ iterations with a reduced learning rate of $\alpha_{1}<\alpha_{0}$, and so on, until the engineer is satisfied with the solution.
-Essentially, the reducing step-size regime involves chaining together a finite number (say $T$) of constant step-size GD runs, with the $(i+1)^{th}$ run having an initialization of $\beta_{u}$ and a learning rate of $\alpha_{i}$, where $u=\sum_{j=0}^{i-1}N_{j}$.
-In the reducing step size regime, the theoretical guarantees described in the diminishing step-size regime may not be realized.
-Nevertheless, the limiting dynamics of the sequence in this regime are exactly those of a constant step-size GD, with learning rates and weight initializations chosen corresponding to the values at the beginning of the $T^{th}$ run.
-Accordingly, in this paper, we exclusively study the constant step-size problem, assuming that the weights and learning rates have been appropriately chosen.
+In the reducing step-size regime, training involves chaining together a finite number (say $T$) of constant step-size GD runs, with the $(i+1)^{th}$ run having an initialization of $\beta_{u}$ and a learning rate of $\alpha_{i}$, where $u=\sum_{j=0}^{i-1}N_{j}$.
+While the theoretical guarantees of the diminishing step-size regime may not be realized, the limiting dynamics of the sequence in this regime are exactly those of a constant step-size GD, with learning rates and weight initializations chosen corresponding to the values at the beginning of the $T^{th}$ run.
+Accordingly, we exclusively study the constant step-size problem, assuming appropriate weights and learning rates have been chosen.
 2 NDGM and convergence analysis
-Recent theoretical work by Davis et al.
-(2020) has shown that in the diminishing step-size regime, NDGMs converge to a local optimal solution in the theoretical limit under fairly general conditions.
-Unfortunately, these regularity conditions are rarely satisfied in practice.
-To illustrate this point, consider the Assumption A.3 in Davis et al.
-(2020), which requires that in the diminishing step size regime, the sum of the learning rates should diverge.
+Recent theoretical work by (Davis et al., 2020) has shown that in the diminishing step-size regime, NDGMs converge to a local optimal solution in the theoretical limit under fairly general conditions.
+Unfortunately, the conditions of the diminishing step size regime are rarely satisfied in practice.
+To illustrate this point, consider the Assumption A.3 in (Davis et al., 2020), which requires that in the diminishing step size regime, the sum of the learning rates should diverge.
 As was previously stated, this is a standard assumption even in the thoeretical convergence analysis of GDs.
@@
 As seen in Figure 1, although the loss function is a monotone decreasing function of the iteration number, the infinity norm of the NDGM sequence eventually takes values greater than 1, hence establishing the contradiction.
-The non-capturing described in Figure 1 is to be expected; as stated in Section 1 of Boyd et al.
-(2003), ”Unlike the ordinary gradient method, the subgradient method is not a descent method; the function value can (and often does) increase”.
+The non-capturing described in Figure 1 is to be expected; as stated in Section 1 of (Boyd et al., 2003), ”Unlike the ordinary gradient method, the subgradient method is not a descent method; the function value can (and often does) increase”.
 Figure 1: Plot of the loss function ($L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$), and infinity norm ($||\beta_{k}||_{\infty}$) of the parameter vector as a function of the iteration number ($k$).
@@
 Given that the dynamics of GDs and NDGMs differ, the question arises what assumptions are appropriate for the study of convergence analysis in neural network training.
-Majority of the analyses in the literature (Défossez et al.
-(2020); Reddi et al.
-(2019); You et al.
-(2019); Zhang et al.
-(2022) to state a few) assume that the loss function, $f(\beta)$ is continuously differentiable, and is $L-smooth$, i.e., there exists a constant $L$, such that
+Majority of the analyses in the literature ((Bottou et al., 2018; Zhang et al., 2022; Défossez et al., 2020; You et al., 2019) to state a few) assume that the loss function, $f(\beta)$ is continuously differentiable, and is $L-smooth$, i.e., there exists a constant $L$, such that
 $||\nabla f(x)-\nabla f(y)||\leq L||x-y||$ (5)
@@
 However, this assumption breaks down for non-differentiable functions, where the gradient does not even exist at various points in the domain.
-This leads us to a pertinent inquiry: Can we salvage this assumption by confining its validity to a local context, possibly within a specific interval centered around the local optimal solution?
+This raises the question if the assumption can be salvaged by confining its validity to a local context around the local optimal solution.
 Unfortunately, even this cannot be done; as stated in (Laurent & Brecht, 2018), “local minima of ReLU networks are generically nondifferentiable.
@@
 $||g(x)-g(y)||\leq L||x-y||$ (6)
-would take $O(1/\epsilon^{2})$ steps (see Section 7.1.3 of Tibshirani (2015) for the result, and a visual illustration of how much slower $O(1/\epsilon^{2})$ is than $O(1/\epsilon)$).
-Thus, it is likely that the vast literature on the convergence analysis of GDs for loss functions satisfying (5) likely have little bearing on the dynamics of convergence in modern non-smooth neural networks.
+would take $O(1/\epsilon^{2})$ steps (see Section 7.1.3 of (Tibshirani, 2015) for the result, and a visual illustration of how much slower $O(1/\epsilon^{2})$ is than $O(1/\epsilon)$).
+Thus, it is likely that the vast literature on the convergence analysis of GDs for loss functions satisfying (5) have little bearing on the dynamics of convergence in modern non-smooth neural networks.
 More generally, we expect the speed of convergence for smooth neural networks to be faster than that of non-smooth neural networks.
 Experiments with NDGM convergence:
-To validate our intuition in a more general context, we conduct experiments on CNNs of various depths (see the appendix for architecture details).
-For each network, we carry out two identical simulations, with the only difference being the activation function used.
-In the first simulation, we employ the non-smooth ReLU activation function, and in the second, we use the similarly shaped, but smooth and asymptotally equivalent, GELU activation function (Hendrycks & Gimpel, 2016).
-As seen in Figure 2, the difference in convergence speeds between smooth and non-smooth networks widens as the depth of the neural network.
-This observation aligns with our expectations, given that increased depth introduces greater non-smoothness into the system.
-Hendrycks & Gimpel (2016) substantiate a similar trend of accelerated convergence in GELU networks compared to ReLU networks across diverse architectures and datasets (see Figures 2 through 7 in their paper).
-Figure 2: Training curves comparing smooth and non-smooth CNNs (from left to right) of depth 11, 13, and 19 respectively.
-As the network depth increases, the non-smoothness in the system also increases.
-Consequently, the difference in convergence speeds between smooth and non-smooth networks widens with the increasing non-smoothness in the system.
-Do results derived in the L-Smooth setting describe the behavior of all differentiable neural networks?:
+We conduct two sets of experiments on CNNs of varying depths to validate our intuition in a more general context (see the Appendix A for architecture details).
+In each set, we perform two identical simulations for every network, with the only difference being the choice of activation function: one simulation uses the the non-smooth ReLU activation function, while the other uses the smooth and asymptotically equivalent GELU activation function (Hendrycks & Gimpel, 2016).
+The first set of experiments excludes batch normalization, while the second includes it.
+Figure 2 shows that without batch normalization, the gap in convergence speeds between smooth and non-smooth networks widens as network depth increases, aligning with our expectations.
+This trend is further substantiated by (Hendrycks & Gimpel, 2016), who demonstrate accelerated convergence in GELU networks compared to ReLU networks across diverse architectures and datasets (see Figures 2 through 7 in their paper).
+However, with batch normalization, the difference between ReLU and GELU training curves becomes indistinguishable, emphasizing its importance in the training process.
+We conclude that the choice of activation function is the primary factor contributing to the difference in convergence speeds between ReLU and GELU curves with increasing depth, as no other principled reason appears evident.
+Figure 2: Training curves comparing smooth and non-smooth CNNs (from left to right) of depth 11, 13, and 19, respectively.
+The top row shows validation losses without batch normalization, while the bottom row includes batch normalization.
+L-smoothness and the geometry of ReLU networks:
+As stated in (Bottou, 2010), the assumption of $L-smoothness$ ensures that the gradient of the loss function does not change too rapidly with respect to the parameter vector.
+However, this assumption which does not hold for ReLU networks.
+To illustrate this, consider the 1-D case with $q(x)=\text{max}(0,x)$.
+In this case, $\lim_{h\rightarrow 0^{+}}(1/2h)(\nabla q(h)-\nabla q(-h))=\infty$, indicating that the gradient changes arbitrarily quickly at the origin.
+This observation necessitates the use of non-smooth analysis for studying the loss functions of ReLU networks, as detailed in (Laurent & Brecht, 2018).
+The key intuition about non-differentiable functions is that they have kinks corresponding to non-differentiable points, and local minima frequently lie at these kinks.
+As a result, gradient descents which rely on smoothness assumptions cannot access these points of interest.
+This realization serves as one of the primary motivations for the development and application of second-order cone programs for convex non-smooth problems (Boyd et al., 2004)  [^3] See 26:30 to 29:00 in https://rb.gy/ab0aeh.
+The inability of gradient descent methods to effectively maneuver non-smooth functions and the associated pathologies form a recurring theme throughout the remaining sections of this paper.
 In a subsequent section on the Edge of Stability, we show that insights based on the assumption of $L-smoothness$ are not universally applicable even across all Lipschitz continuous, convex, differentiable neural networks.
-We point out that $L-smoothness$ is one of the strongest assumptions which can be applied to a function after convexity.
-This assumption guarantees the existence of the first and second derivatives, and the boundedness of the second derivative—a collection of criteria that are unmet by the majority of the loss functions encountered in the neural network literature.
+$L-smoothness$ is one of the strongest assumptions after convexity, guaranteeing the existence and boundedness of the first and second derivatives—criteria unmet by most loss functions in neural network literature.
+Limitations of our work:
+A common limitation in the current and subsequent sections is that we do not analyze the dynamics of the neural network training process as it is often conducted in practice.
+This is necessitated by the fact that neural networks often involve non-differentiable layers like dropout and max pooling, which are beyond the scope of our study.
+We do not reiterate this limitation in every subsequent section.
 The way forward:
 In this section, we show that NDGMs are slower to converge than GDs.
-Fortunately, most recent explorations into image and language models primarily leverage the transformer architecture (Vaswani et al., 2017) which employs the differentiable GELU activation function (Hendrycks & Gimpel, 2016).
+Fortunately, most recent explorations into image and language models primarily leverage the transformer architecture (Vaswani et al., 2017) with the differentiable GELU activation function (Hendrycks & Gimpel, 2016).
 As a consequence, future iterations of neural networks may predominantly employ pure GDs.
@@
 3 NDGM and the LASSO penalty
-Zero weights play a critical role in model compression (Blalock et al., 2020; Han et al., 2015a, b; Li et al., 2016), and it is desirable to force weights down to zero whenever we can do so without sacrificing too much accuracy.
-A common approach to induce sparsity is to penalize the $L_{1}$ norm of the model weights, with the penalizing factor commonly referred to as the LASSO penalty.
-In the statistics literature, loss functions involving penalized $L_{1}$ norms are typically minimized using specialized algorithms designed for different use cases (Efron et al., 2004; Friedman et al., 2008; Mazumder & Hastie, 2012; Tibshirani, 1996; Tibshirani et al., 2005; Zou & Hastie, 2005).
-However when training neural networks, the penalized $L_{1}$ norm is used with arbitrary non-convex functions, with the understanding that running NDGM on the problem results in a near sparse optimal solution (Bengio, 2012; Goodfellow et al., 2016).
-As stated in Scardapane et al.
-(2017), ”Even if (the LASSO) has a non differentiable point in 0, in practice this rarely causes problems to standard first- order optimizers.
-In fact, it is common to simultaneously impose both weight-level sparsity with the $L_{1}$ norm, and weight minimization using the $L_{2}$ norm, resulting in the so-called elastic net penalization”.
-This expectation is not being met in reality, and there is strong evidence of users struggling with the interpretation of their LASSO results in the PyTorch forums  [^3] https://rb.gy/kh0qr  [^4] https://rb.gy/2vsry  [^5] https://rb.gy/6dfli.
-Given the wide variety of critical applications in which the LASSO finds use (Ghosh et al.
-(2021); Li et al.
-(2011); Li & Sillanpää (2012); Ogutu et al.
-(2012) to state a few), it is critical to understand the reliability of the NDGM solution for the LASSO problem.
+Zero weights are crucial for model compression (Blalock et al., 2020; Han et al., 2015a, b; Li et al., 2016), and inducing sparsity by penalizing the $L_{1}$ norm of model weights is a common approach.
+While the statistics literature employs specialized algorithms for problems involving penalized $L_{1}$ norms (Efron et al., 2004; Friedman et al., 2008; Mazumder & Hastie, 2012; Tibshirani, 1996; Tibshirani et al., 2005; Zou & Hastie, 2005), neural networks often use the penalized $L_{1}$ norm with arbitrary non-convex functions, assuming that NDGM results in a near-sparse optimal solution (Bengio, 2012; Goodfellow et al., 2016b; Scardapane et al., 2017).
+This expectation is not always met, as evidenced by user struggles in PyTorch forums [^4] https://rb.gy/kh0qr [^5] https://rb.gy/2vsry [^6] https://rb.gy/6dfli.
+Given the LASSO’s critical applications (Ghosh et al., 2021; Li et al., 2011; Li & Sillanpää, 2012; Ogutu et al., 2012), understanding the reliability of the NDGM solution for the LASSO problem is essential.
 NDGM solutions for the LASSO problem are not reliable:
@@
 Proposition 3.1.
+Let $\{\beta_{t}\}_{t=0}^{\infty}$ denote the sequence generated by running NDGM on (7), with a constant learning rate, $\alpha$, and the entries in $\beta_{0}$ being uniformly sampled from $[-1,1]$.
+Then $\beta_{k}$ will have all non-zero entries with probability 1 for all $k>0$.
+Furthermore, there exists an integer $N_{0}$, such that for every $N\geq N_{0}$, we can write
+$\beta_{n_{k}+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even},\end{cases}$ (8)
+where $\gamma_{k}\in(0,\alpha\lambda_{1})$ for all $10\leq k\leq P-$ 1.
+Proof.
+See Appendix B for details ∎
+Equation (8) suggests that with the same learning rate and initialization, a larger $\lambda_{1}$ can lead to a NDGM solution with a larger $L_{1}$ norm, defeating the purpose of the LASSO penalty.
+To illustrate this, we note that in the 2D case, with weights initialized as $\beta_{0}=[0.5053,0.5053]$, $\alpha=0.01$, and $\lambda_{1}=1$, the NDGM sequence oscillates between $[0.0053,0.0053]$ and $[-0.0047,-0.0047]$.
+However, with $\lambda_{1}=100$, and the same values of $\beta_{0}$ and $\alpha_{1}$, the solution oscillates between $[0.5053,0.5053]$ and $[-0.4947,-0.4947]$.
+Therefore at convergence, the $L_{1}$ norm is at most 0.0106 for $\lambda_{1}=1$, but at least 0.9894 for $\lambda_{1}=100$.
+Furthermore, reducing the step-size to $\alpha=0.001$ after apparent convergence does not resolve this paradox.
+With $\lambda_{1}=1$, the sequence oscillates between $[0.0003,0.0003]$ and $[-0.0007,-0.0007]$, while with $\lambda_{1}=100$, it oscillates between $[0.0053,0.0053]$ and $[-0.0947,-0.0947]$.
+Appendix C shows that this paradox holds for NDGM solutions of the standard $L_{1}$ regularized linear problem.
+The subgradient method and step-size choices:
+The counter-intuitive behavior seen in the example above can be explained using the subgradient method.
+Since the loss function described in equation (7) is Lipschitz continuous (equation (6)) with $L=\lambda_{1}$, we have from (Boyd et al., 2003) (Section 2 on the Constant Step Size) that
+$\lim_{k\rightarrow\infty}||\beta_{k}||_{1}<\alpha\lambda_{1}.$ (9)
+Therefore, for the same value of $\alpha$, the larger value of $\lambda_{1}$ will result in a larger error in estimation, which is precisely what we are seeing above.
+This illustration also highlights the contrast between the diminishing and reducing step size regimes outlined in the preliminaries.
+Under the diminishing step-size regime, $\alpha\rightarrow 0$ in the theoretical limit, and therefore, Proposition 3.1 guarantees that the optimal solution for (7) will be 0 for all $\lambda_{1}>0$.
+On the other hand, in the reducing step-size regime, $\alpha>0$ at the end of training, and therefore, the error in estimation becomes a critical determinant of the optimal value.
+Connections to other NDGM variants:
+Conventional wisdom in deep learning theory suggests that different NDGM variants, such as NDGM with momentum and RMSProp, converge at varying rates but eventually reach similar near-optimal values (Figure 8.5 of (Goodfellow et al., 2016a), slides 64 onwards in https://rb.gy/ojzpgm  [^7] These slides are from one of the most widely used courses for self teaching; see 1:01:00 - 1:03:00 of https://rb.gy/kaj976 where this precise point is made).
+However, we prove in Appendices D and E that these claims are incorrect even for the toy LASSO problem described in (7).
+Furthermore, our analysis reveals two interesting facts.
+First, that the RMSProp sequence for (7) is agnostic of the value of $\lambda_{1}$ used in the analysis.
+Second, there is a phase transition between the RMSProp and vanilla NDGM training curves characterized by a change in the relative behavior of the training curves of the two algorithms as $\lambda_{1}$ is increased from a value which is much smaller than 1 to a value much larger than 1, which cannot be explained without accounting for the non-differentiability of the $L_{1}$ regularizer.
+These findings highlight the importance of incorporating the non-differentiability of regularizers like the $L_{1}$ norm when analyzing the behavior of NDGM variants.
+Failing to account for non-differentiability results in an incomplete understanding of the optimization dynamics.
+Consequently, it remains unclear how the convergence analysis of algorithms like Adam and RMSProp for differentiable loss functions, as often performed in studies such as (Reddi et al., 2019; Défossez et al., 2020; Zhang et al., 2022), relates to the dynamics of these algorithms’ convergence in non-differentiable settings involving regularizers like the $L_{1}$ norm or activation functions like the ReLU.
+Experiments with deep networks:
+In their seminal work on network pruning, (Han et al., 2015b) propose using $L_{1}$ regularization to encourage weights to approach 0, facilitating the pruning process.
+This approach is widely adopted in the literature (Wen et al., 2016; Xiao et al., 2019; Yang et al., 2019) and industry [^8] See the recommendation of the regularizer in the Training config section of https://rb.gy/20y13m and https://rb.gy/i1sb6s.
+To challenge this suggestion’s general applicability, we conduct two sets of experiments finetuning a pre-trained VGG16 model on CIFAR-10 using $L_{1}$ penalized cross-entropy loss with a constant learning rate of $\alpha=0.003$ for 200 epochs.
+We use $\lambda_{1}=0.0002$ in the first set and $\lambda_{1}=2$ in the second, running three parallel trials within each set using RMSProp, SGD, and Adam.
+At the end of our runs, we make two key observations (Figure 5): 1) The $L_{1}$ norms at convergence differ vastly across the three variants, and 2) for SGD and Adam, larger values of $\lambda_{1}$ result in a larger $L_{1}$ norm at convergence.
+These findings suggest that the relationship between $L_{1}$ regularization and prunable parameters may not be as straightforward as previously thought.
+Indeed, our results show that a larger $\lambda_{1}$ does not necessarily correspond to more prunable parameters (Table 1): with $\lambda_{1}=2$, less than 0.2% of the parameters at convergence with vanilla NDGM have absolute value less than $10^{-5}$, but with $\lambda_{1}=0.002$, 99.98% do.
+Soft-thresholding and ISTA:
+The observation that higher $\lambda_{1}$ values do not yield more prunable parameters indicates that while soft thresholding produces sparse solutions for NDGMs, it does not find optimal solutions.
+This issue extends beyond NDGM to general neural network problems, where any soft thresholding solution yields sparsity but not optimality.
+The root cause is that soft thresholding implicitly solves a generalized ISTA problem (Chen et al., 2023), which assumes the loss function is $L-smooth$ (point 1, Section 3.2 of (Chen et al., 2023)) - an assumption that fails to capture the non-convex and non-differentiable dynamics in general neural network settings, as discussed previously.
+Consequently, applying the generalized ISTA formulation relying on $L-smoothness$ to neural networks necessarily leads to suboptimal solutions.
+NDGM and the way forward with sparsity:
+While NDGMs do not inherently generate sparse solutions, iterative methods have been developed for $L_{1}$ regularized problems (Ma et al., 2019; Siegel et al., 2020).
+The realization that NDGMs lack this capability predates the widespread adoption of deep learning (Xiao, 2009).
+Despite this, a misconception persists in the deep learning literature, suggesting that such algorithms yield sparse solutions comparable to those from traditional machine learning frameworks like glmnet (Friedman et al., 2021) or scikit-learn (Kramer & Kramer, 2016).
+However, there is a fundamental disparity between these frameworks: glmnet focuses on a narrow set of problems, guaranteeing optimality, while PyTorch (Paszke et al., 2017) provides flexibility in model creation, but lacks similar guarantees.
+Attaining a solution at a non-differentiable point is challenging, emphasizing the importance of exploring related studies and tools rather than relying on simplistic implementations when precision is crucial.
+Proposition 3.1 shows that the NDGM sequence for the non-differentiable, Lipschitz continuous, convex loss function described in (7) will not diverge to $\infty$ for any finite value of $\alpha$.
+This is a special case of unstable convergence, a topic we cover in greater detail in the next section.
+4 NDGM and the Edge of Stability
+In recent years, there has been a growing interest on the topic of “unstable convergence”.
+The literature on the subject (Ahn et al., 2022b, a; Arora et al., 2022; Chen & Bruna, 2022; Cohen et al., 2021; Li et al., 2022) is motivated by the finding that unlike convex quadratic forms, gradient descent on neural networks do not diverge to $\infty$ even when the learning rate is greater than $\alpha^{*}=2/\eta$, where $\eta$ is the dominant eigenvalue of the loss function’s Hessian.
+Instead, the loss function has been empirically demonstrated to converge “unstably”, reducing non-montonically on the long run.
+$\alpha=\alpha^{*}$ is referred to in the literature as the ”Edge of Stability”, since it separates regions of ”stable convergence” ($\alpha<\alpha^{*}$) from regions of “unstable convergence” [^9] We define unstable convergence to mean “non-divergence to $\infty$”.
+Our definition is adapted from Definition 1.1 of (Arora et al., 2022).
+However, we deviate from their definition due to its dependence on the existence of the Hessian at all points between the current and next iterate, a condition that need not be satisfied in non-smooth neural networks.
+($\alpha>\alpha^{*}$).
+The derivation of the Edge of Stability condition assumes that the loss function, $f$, is $L-smooth$ (see equation (5)).
+It is conjectured that neural networks demonstrate unstable convergence because the phenomenon is inconsistent with widespread presumptions in the field of optimization (Cohen et al., 2021; Ahn et al., 2024).
+This conjecture is incorrect for Lipschitz continuous, differentiable, convex functions satisfying (6) as shown in the proposition below.
+Proposition 4.1.
+All convex non-smooth loss functions having bounded subgradients, or satisfying equation (6) will show unstable convergence.
+Proof.
+Note that any loss function satisfying (6) has subgradients bounded above by $L$.
+Additionally, as outlined in Section 1 of Boyd et al.
+(2003), the deductions derived from the subgradient method hold equivalently for differentiable functions.
+This equivalence arises from the fact that the only admissible value for the subgradient of a differentiable convex function at any given point is, indeed, its gradient.
+Accordingly, using the convergence properties of the subgradient methods with constant step sizes from Section 2 of Boyd et al.
+(2003), we have that $\lim_{k\rightarrow\infty}f(x_{k})-f^{*}\leq\alpha L^{2}<\infty$, hence the result.
+∎
+Figure 3: Training curves for $\alpha=0.1$, $\alpha=1$, and $\alpha=10$ with the Lipschitz continuous Huber loss function.
+As the learning rate increases, the loss starts oscillating with larger frequencies, but never diverges to $\infty$.
+This is to be expected from Proposition 4.1, since the Huber loss function is convex, non-smooth and has bounded subgradients.
+The toy LASSO problem studied in (7), for which we analytically demonstrated unstable convergence in the previous section, is an example of a non-differentiable convex loss function satisfying Proposition 4.1.
+An example of a differentiable non-smooth convex function satisfying Proposition 4.1 is the Huber loss function, which finds use in many state of the art object detection models (Girshick, 2015; Liu et al., 2016; Ren et al., 2015).
+The Huber loss function is non-smooth because it is once but not twice differentiable.
+For a regression problem involving an arbitrary $50\times 1$ response vector, $\mathbf{y}=[y_{1},y_{2}\ldots y_{N}]^{T}$, and an arbitrary $50\times 200$ data matrix, $\mathbf{Z}$, the Huber loss function is defined as
+$L_{7}(\beta)$ $=\frac{1}{N}\sum l(i),\quad\text{where}$ (10)
+$l(i)$ =$=\begin{cases}\frac{1}{2}(y_{i}-z_{i}^{T}\beta)^{2}&\text{if }\left|(y_{i}-z_{i}^{T}\beta)\right|<1\\ \left(|y_{i}-z_{i}^{T}\beta|-\frac{1}{2}\right)&\text{otherwise}\end{cases}$
+where $z_{i}$ is the $i^{th}$ column of $\mathbf{Z}$.
+Figure 3 shows that the NDGM sequence for (10) does not diverge toward $\infty$, even with a high learning rate, such as $\alpha=10$, which aligns with our expectations.
+Proposition 4.1 suggest that it is unclear and the examples that follow show that the common approach of extrapolating results developed for $L-smooth$ loss functions to the non-smooth or non-differentiable case is problematic.
+To emphasize this point, we provide an example of such an extrapolation in the literature which breaks down in the non-smooth case.
+Subquadratic losses and unstable convergence:
+In section 3.1 of (Ma et al., 2022), the authors use second-order finite differences to analyze the curvature of VGG16 and ResNet near the local minimum, concluding that subquadratic growth explains the edge of stability phenomenon.
+However, their analysis is flawed because ReLU networks have generically non-differentiable local minima (Laurent & Brecht, 2018), where their approximation breaks down.
+The Huber loss function described in equation (10) challenges their statement’s broad applicability.
+This overparameterized convex function has a global minimum of 0, exhibiting quadratic growth nearby.
+(Ma et al., 2022)’s result suggests stable convergence, but Proposition 4 guarantees unstable convergence (Figure 3).
+The discrepancy arises because the Huber loss function lacks the twice differentiability required by Definition 4.1 in (Ma et al., 2022).
+In contrast, $L-smooth$ functions, which are twice differentiable by definition, exhibit subquadratic growth near their local minimum during unstable convergence.
+This example highlights the challenges of extending results developed under strong assumptions like $L-smoothness$ to the general case.
+The way forward:
+We demonstrate that our understanding of the Edge of Stability will be incomplete until we more precisely quantify conditions under which we expect it to occur.
+Furthermore, considering ReLU networks’ non-differentiable minima, without further research into NDGM behavior around non-differentiable points our grasp of training dynamics lacks completeness.
+Consequently, we propose employing the convex non-differentiable single-layer neural network presented in (Kumar, 2023) as the toy problem for theory development.
+Studying this neural network offers several advantages: 1) its convex nature renders it more amenable to analysis compared to non-convex functions, 2) we possess analytical knowledge of the global optimum, and 3) the global optimum exists at a point of non-differentiability—the setup which is particularly relevant to ReLU networks.
+5 Conclusion
+In this paper, we highlight how non-differentiability alters training dynamics across three domains: convergence analysis, the LASSO problem, and the Edge of Stability.
+In each case, we show that problems arise because of incorrect assumptions made about the contour of the loss surface.
+Our results suggest the critical need to reach consensus on a set of assumptions which can be made about the loss surface before progress can be made.
+6 Acknowledgements
+We would like to thank Wenyun Zuo, Boris Ginsburg, and several anonymous reviewers for inputs on earlier drafts of this paper.
+We would also like to thank Shripad Tuljapurkar and Marc Feldman for how to best stare at dynamical systems.
+Appendix A Appendix A - Architecture and code for Figure 2
+The code to construct our plots in Figure 2, was taken from https://github.com/kuangliu/pytorch-cifar/blob/master/models/vgg.py, with three minor modifications.
+First, we removed the batch normalization as part of each layer.
+We did this because, while it is known that batch normalization alters the training dynamics (Ioffe & Szegedy, 2015), and our work is interested in the effect of network smoothness on convergence speed in the absence of confounders.
+Second, we replaced max pool with average pool to ensure differentiability of the layer, and third, we used a simpler layer configuration than one provided in the link.
+The configuration we used is provided below:
+cfg = {
+    ’VGG11’: [16, ’M’, 32, ’M’, 32, 32, ’M’, 64, 64, ’M’, 64, 64, ’M’],
+    ’VGG13’: [16, 16, ’M’, 32, 32, ’M’, 32, 32, ’M’, 64, 64, ’M’, 64, 64, ’M’],
+    ’VGG19’: [16, 16, ’M’, 32, 32, ’M’,
+              32, 32, 32, 32, ’M’,
+              64, 64, 64, 64, ’M’,
+              64, 64, 64, 64, ’M’],
+Appendix B Proof of proposition 3.1
+Proposition 3.1 follows from the two propositions described below.
+Proposition B.1.
 Let $\{\beta_{t}\}_{t=0}^{\infty}$ denote the sequence generated by running NDGM on (7), with a constant learning rate, $\alpha$, and the entries in $\beta_{0}$ being uniformly sampled from $[-1,1]$.
@@
 In the next proposition, we show that the sequence does not converge, but oscillates between two fixed points.
-Proposition 3.2.
-The sequence $\{\beta_{t}\}_{t=0}^{\infty}$ described in the previous proposition does not converge.
-Furthermore, there is an integer $N_{0}$ such that for every $N\geq N_{0}$, the sequence oscillates between two vectors : $\gamma_{1}$ and $\gamma_{2}$, where $||\gamma_{1}||_{\infty}<\alpha\lambda_{1}$, and $||\gamma_{2}||_{\infty}<\alpha\lambda_{1}$.
+Proposition B.2.
+The sequence $\{\beta_{t}\}_{t=0}^{\infty}$ described in the previous proposition does not converge. Furthermore, there exists an integer $N_{0}$, such that for every $N\geq N_{0}$, we can write
+$\beta_{n_{k}+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even},\end{cases}$ (13)
+for some $\gamma_{k}\in(0,\alpha\lambda_{1})$ for all $10\leq k\leq P-$ 1.
 Proof.
 We prove the results assuming $\beta_{0}[k]>0$; the proof is similar when $\beta_{0}[k]<0$.
-From Proposition 3.1, we know that for any $N>0$, $\beta_{N}[k]\neq 0$ with probability 1.
+From Proposition B.1, we know that for any $N>0$, $\beta_{N}[k]\neq 0$ with probability 1.
 Therefore, (11) implies that starting from any $\beta_{0}[k]>0$, the sequence will decrease monotonically till it reaches a value between 00 and $\alpha\lambda_{1}$.
@@
 The result follows by choosing $N_{0}=\text{max}(n_{0},n_{1}\ldots n_{P-1})$. ∎
-Equation (10) suggests the surprising result that with the same learning rate and initialization, a larger value of $\lambda_{1}$ can result in a NDGM solution with a larger $L_{1}$ norm, which completely defeats the purpose of the LASSO penalty.
-To better understand the issue, consider the 2D case, with the weights initialized as $\beta_{0}=[0.5053,0.5053]$.
-With $\lambda_{1}=1$ and $\alpha=0.01$, Proposition 3.2 states that the NDGM sequence will oscillate between $[0.0053,0.0053]$ and $[-0.0047,-0.0047]$, but with $\lambda_{1}=100$, and the same learning rate and starting point, the solution will oscillate between $[0.5053,0.5053]$ and $[-0.4947,-0.4947]$.
-Thus, at the point when NDGM appears to converge, the solution with $\lambda_{1}=1$ will have an $L_{1}$ norm of at most 0.0106, but solution with $\lambda_{1}=100$ case will have an $L_{1}$ norm of atleast 0.9894.
-Furthermore, note that reducing the step-size does not resolve this paradox.
-If, in both cases, we decrease the learning rate to $\alpha=0.001$ once the algorithm seems to have converged, then with $\lambda_{1}=1$, the NDGD sequence will oscillate between $[0.0003,0.0003]$ and $[-0.0007,-0.0007]$.
-However, with $\lambda_{1}=100$, the NDGM sequence will oscillate between $[0.0053,0.0053]$ and $[-0.0947,-0.0947]$, and the paradox persists.
-To demonstrate the same issue in the broader setting, consider the loss function for the general LASSO problem given by
+Appendix C NDGM unreliability for the general LASSO problem
+In this section, we demonstrate via simulations that the show that the problems demonstrated in our toy LASSO problem hold even for the general setting.
+Consider the loss function for the general LASSO problem given by
 $L_{4}(\beta;{\bf{W}},{\bf{y}},\lambda_{1})=\frac{1}{N}||{\bf{y}}-{\bf{W}}\beta||_{2}^{2}+\lambda_{1}||\beta||_{1},$ (15)
@@
 At the end of the two runs, the optimal solution with $\lambda_{1}=0.1$ has an $L_{1}$ norm of 1.62, and the optimal solution with $\lambda_{1}=10$ has an $L_{1}$ norm of 25.4 i.e., a 1000 fold increase in the value of $\lambda_{1}$ results in a more than 15 fold increase in the $L_{1}$ norm of the optimal solution!
-Connections to the subgradient method:
-The counter-intuitive behavior seen in the example above can be explained using the subgradient method.
-Since the loss function described in equation (7) satisfies the Lipschitz condition (see equation (6)) with $L=\lambda_{1}$, we have from Boyd et al.
-(2003) (Section 2 on the Constant Step Size) that
-$\lim_{k\rightarrow\infty}||\beta_{k}||_{1}<\alpha\lambda_{1}.$ (12)
-Therefore, for the same value of $\alpha$, the larger value of $\lambda_{1}$ will result in a larger error in estimation, which is precisely what we are seeing above.
-Diminishing v/s reducing step sizes:
-This illustration also highlights the contrast between the diminishing and reducing step size regimes outlined in the preliminaries.
-Under the diminishing step-size regime, $\alpha\rightarrow 0$ in the theoretical limit, and therefore, Proposition 3.1 guarantees that the optimal solution for (7) will be 0 for all $\lambda_{1}>0$.
-On the other hand, in the reducing step-size regime, $\alpha>0$ at the end of training, and therefore, the error in estimation becomes a critical determinant of the optimal value.
-Experiments with deep networks:
-In their seminal work on network pruning, Han et al.
-(2015b) propose the utilization of $L_{1}$ regularization to encourage weights to approach 0, facilitating the pruning process.
-This approach is widely adopted in the literature (Wen et al., 2016; Xiao et al., 2019; Yang et al., 2019), and is explicitly endorsed in NVIDIA’s guides for training standard object detection models such as RetinaNet and YOLOv4666See the recommendation of the regularizer in the Training config section of https://rb.gy/20y13m and https://rb.gy/i1sb6s.
-To challenge the general applicability of this suggestion, we fine-tune two pretrained VGG16 models using the $L_{1}$-penalized cross entropy loss with a learning rate of $\alpha=0.01$, with one model using a regularization parameter of $\lambda_{1}=0.001$ and the other with $\lambda_{1}=0.1$.
-At the end of fine-tuning, the $L_{1}$ norm of the weights in the model fine-tuned with $\lambda_{1}=0.1$ is more than 30 times larger (69178.44) than that of the model fine-tuned with $\lambda_{1}=0.001$ (2113.83).
-Furthermore, the latter model has 99.7% of its weights less than $10^{-5}$, whereas the former model has only 0.01% of its weights less than $10^{-5}$, further underscoring the discrepency.
-We conducted experiments with different values of $\alpha$, and the widely used Adam optimizer (Kingma & Ba, 2014), and in each case, we observed the same paradox.
-Our code is provided in the appendices.
-NDGM and the way forward with sparsity:
-While NDGMs do not inherently generate sparse solutions, iterative methods have been developed to produce sparse solutions for the $L_{1}$ regularized linear problem (see Table 1 in Bottou (2010)), and $L_{1}$ regularized neural networks in general (Ma et al., 2019; Siegel et al., 2020).
-The realization that NDGMs lack the capability to produce sparse solutions predates the widespread adoption of deep learning.
-As mentioned in (Xiao, 2009), “… stochastic gradient descent (SGD), has limited capability of exploiting problem structure in solving regularized learning problems”.
-Despite this, a prevalent misconception persists in the deep learning literature, erroneously suggesting that such algorithms yield sparse solutions, and that the solutions generated by deep learning frameworks like PyTorch (Paszke et al., 2017) ought to be comparable to those generated using traditional machines learning frameworks like glmnet (Friedman et al., 2021) or scikit-learn (Kramer & Kramer, 2016).
-We point out that there exits a fundamental disparity between traditional machine learning frameworks like glmnet and deep learning frameworks like PyTorch.
-glmnet focuses on a narrow set of optimization problems, concealing implementation intricacies while guaranteeing optimality.
-Conversely, PyTorch provides greater flexibility in constructing models but lacks similar guarantees, placing the responsibility for solution quality on users.
-Attaining a solution at a non-differentiable point poses a challenging problem, underscoring the significance of exploring related studies and tools rather than opting for a simplistic implementation, particularly when precision is paramount.
-Connections to the Edge of Stability:
-Proposition 3.2 shows that the NDGM sequence for the non-differentiable, Lipschitz continuous, convex loss function described in (7) will not diverge to $\infty$ for any finite value of $\alpha$.
-This is a special case of unstable convergence, a topic we cover in greater detail in the next section.
-4 NDGM and the Edge of Stability
-In recent years, there has been a growing interest on the topic of “unstable convergence”.
-The literature on the subject (Ahn et al., 2022a, b; Arora et al., 2022; Chen & Bruna, 2022; Cohen et al., 2021; Li et al., 2022) is motivated by the finding that unlike convex quadratic forms, gradient descent on neural networks do not diverge to $\infty$ even when the learning rate is greater than $\alpha^{*}=2/\eta$, where $\eta$ is the dominant eigenvalue of the loss function’s Hessian.
-Instead, the loss function has been empirically demonstrated to converge “unstably”, reducing non-montonically on the long run.
-$\alpha=\alpha^{*}$ is referred to in the literature as the ”Edge of Stability”, since it separates regions of ”stable convergence” ($\alpha<\alpha^{*}$) from regions of “unstable convergence” [^7] We define unstable convergence to mean “non-divergence to $\infty$”.
-Our definition is adapted from Definition 1.1 of Arora et al.
-(2022).
-However, we deviate from their definition due to its dependence on the existence of the Hessian at all points between the current and next iterate, a condition that need not be satisfied in non-smooth neural networks.
-($\alpha>\alpha^{*}$).
-The derivation of the Edge of Stability condition assumes that the loss function, $f$, is $L-smooth$ (see equation (5)).
-Since the result has been demonstrated in most convex optimization settings (Jacot et al., 2018; Li & Liang, 2018), it is conjectured that the result ought to hold even in the case of non-convex, non-differentiable neural networks.
-This conjecture is incorrect even for Lipschitz continuous, differentiable, convex functions satisfying (6) as shown in the proposition below.
-Proposition 4.1.
-All convex non-smooth loss functions having bounded subgradients, or satisfying equation (6) will show unstable convergence.
+Appendix D Different NGDM variants and the LASSO
+In this section, we demonstrate that various NDGM variants, traditionally thought to converge similarly based on the analysis of differentiable functions, exhibit notably distinct behaviors even when applied to the simple non-differentiable toy LASSO problem outlined in (7).
+A summary of our findings are in Table LABEL:tab:variantcomp.
+NDGM variant Unique characteristic
+Vanilla NDGM (including SGD) $\beta_{0}$ and $\lambda_{1}$ influence eventual values of the sequence. If sequence hits 0, then stays 0.
+NDGM with momentum If sequence hits 0, then jumps away from 0.
+RMSProp Only $\beta_{0}$ influences the eventual value of the sequence. If sequence hits 0, then stays 0.
+D.1 SGD and the LASSO
+Since the gradient of (7) is independent of the batch in question, the results in Proposition B.2 hold for the SGD case as well.
+D.2 NDGM with momentum and the LASSO
+With the vanilla NDGM, (11) gives the guarantee that once a parameter value reaches zero during an iteration, it remains zero in all future iterations.
+We show that NDGM with momentum does not provide this guarantee.
+Proposition D.1.
+Let $\{\beta_{t}\}_{t=0}^{\infty}$ denote the sequence generated by running NDGM with momentum on (7), with a constant learning rate, $\alpha$, and the entries in $\beta_{0}$ being uniformly sampled from $[-1,1]$.
+Suppose $\beta_{N-1}[k]\neq 0$ and $\beta_{N}[k]=0$, then $\beta_{N+1}[k]\neq 0$.
 Proof.
-Note that any loss function satisfying (6) has subgradients bounded above by $L$.
-Additionally, as outlined in Section 1 of Boyd et al.
-(2003), the deductions derived from the subgradient method hold equivalently for differentiable functions.
-This equivalence arises from the fact that the only admissible value for the subgradient of a differentiable convex function at any given point is, indeed, its gradient.
-Accordingly, using the convergence properties of the subgradient methods with constant step sizes from Section 2 of Boyd et al.
-(2003), we have that $\lim_{k\rightarrow\infty}f(x_{k})-f^{*}\leq\alpha L^{2}<\infty$, hence the result.
-∎
-Figure 3: Training curves for $\alpha=0.1$, $\alpha=1$, and $\alpha=10$ with the Lipschitz continuous Huber loss function.
-As the learning rate increases, the loss starts oscillating with larger frequencies, but never diverges to $\infty$.
-This is to be expected from Proposition 4.1, since the Huber loss function is convex, non-smooth and has bounded subgradients.
-In the previous section, we analytically derive the unstable convergence for the LASSO problem described in (7).
-An example of a differentiable non-smooth convex function satisfying Proposition 4.1 is the Huber loss function, which finds use in many state of the art object detection models (Girshick, 2015; Liu et al., 2016; Ren et al., 2015).
-The Huber loss function is non-smooth because it is once but not twice differentiable.
-For a regression problem involving an arbitrary $50\times 1$ response vector, $\mathbf{y}=[y_{1},y_{2}\ldots y_{N}]^{T}$, and an arbitrary $50\times 200$ data matrix, $\mathbf{Z}$, the Huber loss function is defined as
-$L_{7}(\beta)$ $=\frac{1}{N}\sum l(i),\quad\text{where}$ (13)
-$l(i)$ =$=\begin{cases}\frac{1}{2}(y_{i}-z_{i}^{T}\beta)^{2}&\text{if }\left|(y_{i}-z_{i}^{T}\beta)\right|<1\\ \left(|y_{i}-z_{i}^{T}\beta|-\frac{1}{2}\right)&\text{otherwise}\end{cases}$
-where $z_{i}$ is the $i^{th}$ column of $\mathbf{Z}$.
-Figure 3 shows that the NDGM sequence for (13) does not diverge toward $\infty$, even with a high learning rate, such as $\alpha=10$, which aligns with our expectations.
-Furthermore, the figure reveals an intriguing loss pattern when $\alpha=1$: the loss initially increases before the oscillations become noticeable.
-This peculiar behavior has been documented in the literature (Ahn et al., 2022b; Arora et al., 2022), although a conclusive explanation for this phenomenon remains elusive at this time.
-Proposition 4.1 suggest that it is unclear and the examples that follow show that the common approach of extrapolating results developed for $L-smooth$ loss functions to the non-smooth or non-differentiable case is problematic.
-In the next subsection, we demonstrate the problems with this approach by providing an example of an extrapolation which breaks down in the non-smooth case.
-Subquadratic losses and unstable convergence
-In section 3.1 of Ma et al.
-(2022), the authors employ second-order finite differences to investigate the curvature of VGG16 and ResNet in the vicinity of the local minimum.
-They conclude that subquadratic growth of the loss function around this minimum point explains the edge of stability phenomenon.
-However, their analysis is flawed because VGG16 and ResNet, as ReLU networks, have local minima that are generically non-differentiable (Laurent & Brecht, 2018).
-As a result, the points of interest lie where their approximation breaks down.
-To challenge the broad applicability of their statement, consider the Huber loss function depicted in Equation (13).
-This function is overparameterized and therefore possesses a global minimum of 0.
-Consequently, within the vicinity of this global minimum, the loss function exhibits quadratic growth.
-In such instances, Ma et al.
-(2022)’s result states that we will not see unstable convergence.
-However, contrary to their suggestions, Proposition 4 guarantees unstable convergence under these circumstances, as evidenced in Figure 3.
-The discrepancy arises because the Huber loss function lacks the twice differentiability required by Definition 4.1 of Ma et al.
-(2022)’s paper.
-In contrast, all $L-smooth$ functions will exhibit subquadratic growth near their local minimum during unstable convergence, as noted by (Ma et al., 2022).
-This occurs because $L-smooth$ functions are twice differentiable by definition.
-This example highlights the challenges of extending results developed under strong assumptions like $L-smoothness$ to the general case.
-The way forward:
-We demonstrate that our understanding of the Edge of Stability will be incomplete until we more precisely quantify conditions under which we expect it to occur.
-Furthermore, considering ReLU networks’ non-differentiable minima, without further research into NDGM behavior around non-differentiable points our grasp of training dynamics lacks completeness.
-Consequently, we propose employing the convex non-differentiable single-layer neural network presented in (Kumar, 2023) as the toy problem for theory development.
-Studying this neural network offers several advantages: 1) its convex nature renders it more amenable to analysis compared to non-convex functions, 2) we possess analytical knowledge of the global optimum, and 3) the global optimum exists at a point of non-differentiability—the setup which is particularly relevant to ReLU networks.
-5 Conclusion
-In this paper, we highlight how non-differentiability alters training dynamics across three domains: convergence analysis, the LASSO problem, and the Edge of Stability.
-In each case, we show that problems arise because of incorrect assumptions made about the contour of the loss surface.
-Our results suggest the critical need to reach consensus on a set of assumptions which can be made about the loss surface before progress can be made.
-Appendix A Appendix A - Architecture and code for Figure 2
-The code to construct our plots in Figure 2, was taken from https://github.com/kuangliu/pytorch-cifar/blob/master/models/vgg.py, with three minor modifications.
-First, we removed the batch normalization as part of each layer.
-We did this because, while it is known that batch normalization alters the training dynamics (Ioffe & Szegedy, 2015), and our work is interested in the effect of network smoothness on convergence speed in the absence of confounders.
-Second, we replaced max pool with average pool to ensure differentiability of the layer, and third, we used a simpler layer configuration than one provided in the link.
-The configuration we used is provided below:
-cfg = {
-    ’VGG11’: [16, ’M’, 32, ’M’, 32, 32, ’M’, 64, 64, ’M’, 64, 64, ’M’],
-    ’VGG13’: [16, 16, ’M’, 32, 32, ’M’, 32, 32, ’M’, 64, 64, ’M’, 64, 64, ’M’],
-    ’VGG19’: [16, 16, ’M’, 32, 32, ’M’,
-              32, 32, 32, 32, ’M’,
-              64, 64, 64, 64, ’M’,
-              64, 64, 64, 64, ’M’],
-Appendix B Appendix B - Code for paradoxes relating to the LASSO
-The training code was generated using ChatGPT and verified for correctness. The code below utilizes SGD with a learning rate of 0.01, though can trivially be changed to any other.
+With a momentum factor of $\eta$, NGDM with momentum can be written as (see page 6 of https://www.ceremade.dauphine.fr/~waldspurger/tds/22_23_s1/advanced_gradient_descent.pdf for details)
+$\beta_{N+1}[k]=\beta_{N}[k]-\alpha(1-\eta)\nabla L_{3}(\beta_{N})[k]+\eta(\beta_{N}[k]-\beta_{N-1}[k])$ (16)
+Since $\beta_{N}[k]=0$, we have $\nabla L_{3}(\beta_{N}[k])=0$, and therefore,
+$\beta_{N+1}[k]=-\eta\beta_{N-1}[k]\neq 0,$ (17)
+hence the result. ∎
+D.3 RMSProp and the LASSO
+Proposition B.2 shows that the vanilla NDGM sequence eventually bounces between two vectors which depend on the strength of regularization, $\lambda_{1}$, and the weight initialization, $\beta_{0}$.
+Here we show that RMSProp, which is a near scale invariant NDGM sequence bounces between two vectors whose orientation depends only on $\beta_{0}$.
+The RMSProp equations (equation 18 in (Ruder, 2016)) can be written as
+$v_{t}[k]$ $=\gamma v_{t-1}[k]+(1-\gamma)\left(\nabla L_{3}(\beta_{N})[k]\right)^{2},\mbox{ and }$ (18)
+$\beta_{t+1}[k]$ $=\beta_{t}[k]-\frac{\alpha}{\sqrt{v_{t}[k]+\bar{\epsilon}}}\nabla L_{3}(\beta_{t})[k],$ (19)
+where $v_{0}[k]=0$, $\gamma$ is a scaling factor with a typical default value of 0.99, and $\bar{\epsilon}$ is a small positive jitter added to prevent division by 0.
+If $\beta_{N}[k]\neq 0$, we have $\left(\nabla L_{3}(\beta_{N})[k]\right)^{2}=\lambda_{1}^{2}$.
+Using this fact in (18), we note that $v_{1}[k]=(1-\gamma)\lambda_{1}^{2}$, and $v_{2}[k]=(1-\gamma^{2})\lambda_{1}^{2}$.
+Completing the recursion we get
+$v_{t}[k]=(1-\gamma^{t})\lambda_{1}^{2}.$ (20)
+Plugging (20) into (19) with $\epsilon=\bar{\epsilon}/\lambda_{1}^{2}$ gives
+$\beta_{t+1}[k]=\beta_{t}[k]-\frac{\alpha}{\sqrt{1+\epsilon-\gamma^{t}}}sign(\beta_{t}[k]),$ (21)
+which we use as a base in our analysis. We now provide the key results for RMSProp:
+Proposition D.2.
+Let $\{\beta_{t}\}_{t=0}^{\infty}$ denote the sequence generated by running the RMSProp recursion described in (21), with a constant learning rate, $\alpha$, and the entries in $\beta_{0}$ being uniformly sampled from $[-1,1]$.
+Then $\beta_{t}$ will have all non-zero entries with probability 1 for all $t>0$.
+Suppose the the $k^{th}$ entry in the RMSProp sequence vector becomes 0 for the first time after $N$ iterations. Then, completing the recursion in (21) we have
+$\beta_{N}[k]=\beta_{0}[k]+\sum_{i\in Q}\frac{\alpha}{\sqrt{1+\epsilon-\gamma^{i}}}-\sum_{j\in P}\frac{\alpha}{\sqrt{1+\epsilon-\gamma^{j}}}=0,$ (22)
+where $Q$ is the set of all iteration numbers where the parameter value is less than 0 and $P$ is the set of all iteration numbers where the parameter value is greater than 0, with $|P|+|Q|=N$.
+Let $C(N)$ denote the set of all values of $\beta_{0}[k]$ which satisfy (22).
+The cardinality of $C(N)$ is atmost equal to the number of ways of partitioning $N$ integers into 2 sets, $P$ and $Q$.
+Therefore, $C(N)$ is countable, and the superset of all feasible values of $\beta_{0}[k]$ which lead to a sparse solution given by $\bigcup_{I=1}^{\infty}C(I)$ is countable.
+Since the initializations are uncountable, the probability of the event has measure 0.
+Proposition D.3.
+The RMSProp sequence described in (21) eventually behaves like a vanilla NGDM sequence with a LASSO penalty of 1 irrespective of the value of $\lambda_{1}$ chosen in the original problem.
+Proof.
+From the previous proposition, we know that $\beta_{t}[k]\neq 0$ with probability 1 even for large values of $t$.
+Furthermore, we note that even after a modest number of iterations the scaling term becomes negligible; for instance, with $t=1000$, $\gamma^{t}=O(10^{-5})$.
+Accordingly, assuming that in the large $t$ limit we have $\epsilon\ll\gamma^{t}\ll 1$, we can use first order Taylor expansions to write
+$\beta_{t+1}[k]=\beta_{t}[k]-\alpha sign(\beta_{t}[k])+O(\gamma^{t}),$ (23)
+which is the recursion for the vanilla NGDM algorithm with a LASSO penalty of 1 up to first order. ∎
+Appendix E Connections to deep learning training theory:
+The conventional wisdom in deep learning is that different NDGM variants converge at varying rates, but eventually reach similar near-optimal values by the end of training.
+In this section, we demonstrate problems with this assumption using our toy LASSO problem from Equation (7).
+Since the vector of zeros is the unique global minimizer of Equation (7), we expect the vanilla NDGM and RMSProp sequences to converge to this value by the end of training.
+Accordingly, we anticipate the $L_{1}$ norms of the sequences generated using these two variants to converge to comparable near-zero values after training completes.
+Because the loss function in Equation (7) is the $L_{1}$ norm of the weights scaled by a factor of $\lambda_{1}$, we expect the NDGM and RMSProp training curves to also hover around 0 at the end of training, exhibiting substantial overlap.
+To test this hypothesis, we ran vanilla NDGM and RMSProp on Equation (7) with identical initializations and learning rates, plotting the final 100 training iterations for each sequence using two different $\lambda_{1}$ values.
+As Figure 4 shows, the two sequences display no overlap for either $\lambda_{1}$ value.
+Moreover, the RMSProp solution consistently has a higher loss (and $L_{1}$ norm) across all 100 iterations when $\lambda_{1}$ = 0.001, while vanilla NDGM always yields a higher value when $\lambda_{1}$ = 1.
+Conventional analysis of differentiable functions cannot explain this peculiarity, but our findings provide intuition into the underlying reasons.
+To explain these findings, we use the following proposition:
+Figure 4: Comparison of the training curves of RMSProp and vanilla NDGM with $\lambda_{1}=0.001$, and $\lambda_{1}=100$ for the toy LASSO problem considered in (7) with $P=10$ and $\alpha=0.1$.
+We note that when $\lambda_{1}\ll 1$, the RMSProp training curve lies above the vanilla NDGM training curve, and the reverse is true for $\lambda_{1}\gg 1$.
+Proposition E.1.
+For the vanilla NDGM sequence we have $0<\liminf_{N\rightarrow\infty}||\beta_{N}||_{1}\leq\limsup_{N\rightarrow\infty}||\beta_{N}||_{1}<P\alpha\lambda_{1}$, where $P$ is the number of rows in $\beta_{0}.$
+Proof.
+[…]
+We know from propositions B.2 and D.3 that the vanilla NDGM sequence and the RMSProp sequence eventually bounce (approximately) between two points each.
+Let $v_{1}(\beta_{0},\lambda_{1})$ and $v_{2}(\beta_{0},\lambda_{1})$ indicate the points around which the vanilla NDGM sequence bounces, and let $r_{1}(\beta_{0},\lambda_{1})$ and $r_{2}(\beta_{0},\lambda_{1})$ indicate the points around which the RMSProp sequence bounces.
+From propositions E.1 and B.2 we have that $\mbox{min}(\lambda_{1}||v_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||v_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}^{2}$, and $\mbox{max}(\lambda_{1}||v_{1}(\beta_{0},\lambda_{1}||_{1}),\lambda_{1}||v_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}^{2}$.
+From proposition D.3 we have that the RMSProp sequence eventually behaves like a vanilla NDGM sequence with a LASSO penalty of 1, and therefore, we have $\mbox{min}(\lambda_{1}||r_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||r_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}$, and $\mbox{max}(\lambda_{1}||r_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||r_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}$.
+Therefore, when $\lambda_{1}\ll 1$, we intuitively expect $\mbox{max}(\lambda_{1}||v_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||v_{2}(\beta_{0},\lambda_{1})||_{1})<\mbox{min}(\lambda_{1}||r_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||r_{2}(\beta_{0},\lambda_{1})||_{1})$, and when $\lambda_{1}\gg 1$ we expect $\mbox{max}(\lambda_{1}||r_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||r_{2}(\beta_{0},\lambda_{1})||_{1})<\mbox{min}(\lambda_{1}||v_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||v_{2}(\beta_{0},\lambda_{1})||_{1})$, which is what the figure shows.
+To clarify the ideas, suppose $P=10$ and $\alpha=0.1$.
+The above propositions state that with $\lambda_{1}=0.01$, the RMSProp loss oscillates between two positive values, each less than 0.01, while the vanilla NDGM loss oscillates between two positive values, each less than 0.0001.
+Consequently, when $\lambda_{1}$ is significantly less than 1, it is reasonable to anticipate the RMSProp loss will typically be greater than the vanilla NDGM loss at training’s end.
+However, when $\lambda_{1}=100$, the RMSProp loss oscillates between two positive values under 100, whereas the vanilla NDGM loss oscillates between two values under 10,000.
+Accordingly, when $\lambda_{1}$ is significantly greater than 1, the RMSProp loss is expected to be less than the vanilla NDGM loss in most scenarios at training’s conclusion, as demonstrated in the figure.
+Appendix F Experiments with the LASSO penalty:
+In this section, we provide the graphs and tables resulting from the experiments outlined in Section 3, along with the respective code to generate them.
+F.1 $L_{1}$ norm v/s iteration number using different NDGM variants
+Figure 5: Comparison of the training curves of RMSProp, SGD and Adam for a pretrained VGG16 finetuned using the $L_{1}$ penalized cross entropy loss.
+All 6 training runs are initialized with the same weights, use the same learning rate, and are fed in batches in the same sequence.
+F.2 Percentage of weights less than $10^{-5}$ for different NDGM variants
+Variant $\lambda_{1}=0.002$ $\lambda_{1}=2$
+SGD 99.89% 0.18%
+RMSProp 0.09% 1.21%
+Adam 3.31% 2.28%
+Table 1: A comparison of the percentage of weights having absolute value less than $10^{-5}$ across the three NDGM variants.
+We note for Adam and SGD that increasing the $L_{1}$ norm can lead to a reduction in the percentage of prunable parameters.
+F.3 Code for generating the above plots (GPT generated and validated for correctness)
 import torch
@@
 from torchvision import datasets, transforms, models
+import copy
 # Set random seed for reproducibility
@@
     transforms.Normalize((0.5, 0.5, 0.5), (0.5, 0.5, 0.5)),
-train_dataset = datasets.CIFAR10(root=’./data’, train=True, download=True, transform=transform)
-train_loader = torch.utils.data.DataLoader(train_dataset, batch_size=512, shuffle=True, num_workers=4)
-# Load pretrained VGG16 model
-model = models.vgg16(pretrained=True)
-model.to(device)
-# Define L1 regularization strength
-l1_strength = 0.1
-# Define the loss function (CrossEntropyLoss + L1 penalty)
-criterion = nn.CrossEntropyLoss()
-optimizer = optim.SGD(model.parameters(), lr=0.01)
-# Training loop
-epochs = 150
-for epoch in range(epochs):  # Adjust the number of epochs as needed
+train_dataset = datasets.CIFAR10(root=’./data’,
+                                 train=True,
+                                 download=True,
+                                 transform=transform)
+train_loader = torch.utils.data.DataLoader(train_dataset,
+                                           batch_size=512,
+                                           shuffle=False,
+                                           num_workers=4)
+test_dataset = datasets.CIFAR10(root=’./data’,
+                                train=False,
+                                download=True,
+                                transform=transform)
+test_loader = torch.utils.data.DataLoader(test_dataset,
+                                          batch_size=128,
+                                          shuffle=False,
+                                          num_workers=2)
+def compute_validation_loss(net):
+    val_loss = 0.0
+    total = 0
+    with torch.no_grad():
+        for data in test_loader:
+            images, labels = data
+            images, labels = images.to(device), labels.to(device)
+            outputs = net(images)
+            loss = criterion(outputs, labels)
+            val_loss += loss.item()
+            total += labels.size(0)
+    return val_loss / total
+def train(model, optimizer, l1_strength, num_epochs):
+    train_losses, val_losses,l1_norms = [],[],[]
+    for epoch in range(num_epochs):
+        epoch_loss = 0
         for inputs, labels in train_loader:
@@
             optimizer.step()
-    print(f’Epoch [{epoch+1}/100], L1 Penalty: {l1_penalty.item():.4f}’)
-print("Training complete!")
-total_params = 0
-num_small_params = 0
-threshold = 1e-5
+            epoch_loss += loss.item()
+        l1_norms.append(l1_penalty.item())
+        train_losses.append(epoch_loss / len(train_loader))
+        val_losses.append(compute_validation_loss(model))
+        print(f"Epoch: {epoch}, loss: {train_losses[-1]}, val loss: {val_losses[-1]}")
+    return train_losses, val_losses, l1_norms
+# Model initialization
+vgg16 = models.vgg16(pretrained=True)
+initial_state = copy.deepcopy(vgg16.state_dict())
+vgg16.to(device)
+optimizers = {
+    ’SGD’: optim.SGD(vgg16.parameters(), lr=0.003),
+    ’RMSProp’: optim.RMSprop(vgg16.parameters(), lr=0.003),
+    ’Adam’: optim.Adam(vgg16.parameters(), lr=0.003)
+def get_num_prunable_params(model, threshold = 1e-5):
+    total_params, num_small_params = 0,0
     for name, param in model.named_parameters():
@@
             num_small_params += torch.sum(torch.abs(param) < threshold).item()
-fraction_small_params = num_small_params / total_params
-print(f’Fraction of parameters with values less than 10^-5: {fraction_small_params:.5f}’)
+    return num_small_params / total_params
+loss_curves = {}
+l1_norms_curves = {}
+validation_loss_curves = {}
+num_prunable_params = {}
+for name, opt in optimizers.items():
+    loss_curves[name],
+    validation_loss_curves[name],
+    l1_norms_curves[name],
+    num_prunable_params[name] = {},{},{},{}
+    for lambda1 in [2,0.002]:
+        print(f"opt: {name}, lambda: {lambda1}")
+        vgg16.load_state_dict(copy.deepcopy(initial_state))  # Reset model
+        losses, validation_losses, l1_norms = train(vgg16, opt, lambda1, 200)
+        num_prunable_params[name][lambda1] = get_num_prunable_params(vgg16)
+        loss_curves[name][lambda1] = losses
+        validation_loss_curves[name][lambda1] = validation_losses
+        l1_norms_curves[name][lambda1] = l1_norms

### v4

@@
 Abstract
-This paper investigates the distinctions between gradient methods applied to non-differentiable functions (NGDMs) and classical gradient descents (GDs) designed for differentiable functions.
-First, we demonstrate significant differences in the convergence properties of NGDMs compared to GDs, challenging the applicability of the extensive neural network convergence literature based on $L-smoothness$ to non-smooth neural networks.
-Next, we demonstrate the paradoxical nature of NGDM solutions for $L_{1}$-regularized problems, showing that increasing the regularization penalty leads to an increase in the $L_{1}$ norm of optimal solutions in NGDMs.
-Consequently, we show that widely adopted $L_{1}$ penalization-based techniques for network pruning do not yield expected results.
-Additionally, we dispel the common belief that optimization algorithms like Adam and RMSProp perform similarly in non-differentiable contexts.
-Finally, we explore the Edge of Stability phenomenon, indicating its inapplicability even to Lipschitz continuous convex differentiable functions, leaving its relevance to non-convex non-differentiable neural networks inconclusive.
-Our analysis exposes misguided interpretations of NGDMs in widely referenced papers and texts due to an overreliance on strong smoothness assumptions, emphasizing the necessity for a nuanced understanding of foundational assumptions in the analysis of these systems.
-Machine Learning, ICML
-Gradient Descent (GD) and its variants (Duchi et al., 2011; Kingma & Ba, 2014; Lydia & Francis, 2019; McMahan & Streeter, 2010; Shi & Li, 2021; Tieleman et al., 2012; Zhang, 2018) have played a pivotal role in advancing image and language processing over the last two decades.
-These algorithms have found widespread use in neural network training because the are easy to implement, highly scalable, and iteratively approach stationary points, where the loss function’s gradient is zero (Bertsekas, 1997).
-Despite being initially designed for differentiable loss functions, GDs are widely used to minimize the loss functions of non-differentiable neural networks.
-The loss functions of these non-differentiable neural networks takes a special form – they are differentiable almost everywhere.
-Consequently, when gradient descents are applied to these non-differentiable neural networks (which we refer to as non-differentiable gradient methods or NGDMs), they seldom encounter non-differentiable points during the training process.
-This trait, combined with recent findings that NGDMs converge to stationary points under mild regularity conditions (Bolte & Pauwels, 2021; Davis et al., 2020), has led standard references in the field to assume that non-differentiability in neural networks is inconsequential (see Section 6.1.5 of (Stevens et al., 2020), for instance).
-To emphasize this point, we note that three of the most widely cited surveys on optimization methods for neural networks (Bottou et al., 2018; Le et al., 2011; Ruder, 2016) do not mention non-differentiability, even though the ReLU, arguably the most common activation function in neural networks, is non-differentiable.
-Moreover, this assumption has sparked renewed interest in the dynamics of gradient descents amid complex loss landscapes ((Ahn et al., 2022b; Dauphin et al., 2014; Reddi et al., 2019) to state a few), with fewer studies explicitly addressing non-differentiability (Bertoin et al., 2021; Bolte & Pauwels, 2021; Davis et al., 2020).
-Additionally, it has led theoretical papers to develop results for continuously differentiable functions and then claim universality by running simulations on non-differentiable neural networks (Experiment 6 in (Ahn et al., 2022b), Section 3.1 in (Ma et al., 2022), Section 6 in (Zhang et al., 2022) to state a few).
-In our paper, we challenge this assumption, demonstrating that it leads to inaccuracies in several highly referenced papers and texts regarding training dynamics in neural networks.
-We analyze three different aspects of the training process and make the following contributions:
-  • Convergence Analysis of ReLU Networks: We show that the rate of convergence of NGDMs is slower than that of GDs.
-Our findings strongly suggest that the many claims about rate of convergence made under the assumption of $L-smoothness$ have little bearing on the properties of NGDMs.
-Furthermore, we dispel the widespread belief that different NDGM variants like RMSProp and Adam largely differ only in their rates of convergence, but converge to roughly the same eventual point in the limit.
-  • Solution to the LASSO problem: We show that NGDMs, including variants like RMSProp and NDGM with momentum, do not produce sparse solutions for the LASSO problem, even for the simplest case of the $L_{1}$ penalized linear model, and exhibit different behaviors, contrary to prevalent wisdom.
-Moreover, NGDMs can yield counterintuitive results, such as producing a solution with a larger $L_{1}$ norm with a larger LASSO penalty.
+This paper critically examines the fundamental distinctions between gradient methods applied to non-differentiable functions (NGDMs) and classical gradient descents (GDs) for differentiable functions, revealing significant gaps in current deep learning optimization theory.
+We demonstrate that NGDMs exhibit markedly different convergence properties compared to GDs, strongly challenging the applicability of extensive neural network convergence literature based on $L-smoothness$ to non-smooth neural networks.
+Our analysis reveals paradoxical behavior of NDGM solutions for $L_{1}$-regularized problems, where increasing regularization counterintuitively leads to larger $L_{1}$ norms of optimal solutions.
+This finding calls into question widely adopted $L_{1}$ penalization techniques for network pruning.
+We further challenge the common assumption that optimization algorithms like RMSProp behave similarly in differentiable and non-differentiable contexts.
+Expanding on the Edge of Stability phenomenon, we demonstrate its occurrence in a broader class of functions, including Lipschitz continuous convex differentiable functions.
+This finding raises important questions about its relevance and interpretation in non-convex, non-differentiable neural networks, particularly those using ReLU activations.
+Our work identifies critical misunderstandings of NDGMs in influential literature, stemming from an overreliance on strong smoothness assumptions.
+These findings necessitate a reevaluation of optimization dynamics in deep learning, emphasizing the crucial need for more nuanced theoretical foundations in analyzing these complex systems.
+1 Introduction
+Gradient Descent (GD) and its variants (Duchi et al., 2011; Kingma and Ba, 2014; Lydia and Francis, 2019; McMahan and Streeter, 2010; Shi and Li, 2021; Tieleman et al., 2012; Zhang, 2018) have been instrumental in advancing image and language processing over the last two decades.
+These algorithms have gained widespread adoption in neural network training due to their ease of implementation, high scalability, and ability to iteratively approach stationary points where the loss function’s gradient vanishes (Bertsekas, 1997).
+However, an important consideration has emerged in the field: despite being originally designed for differentiable loss functions, GDs are routinely applied to minimize loss functions of non-differentiable neural networks.
+These non-differentiable loss functions possess a unique characteristic—they are differentiable almost everywhere.
+Consequently, when gradient descents are applied to these non-differentiable neural networks (which we term non-differentiable gradient methods or NGDMs), they rarely encounter non-differentiable points during training.
+This property, coupled with recent findings suggesting NGDMs converge to stationary points under mild regularity conditions (Bolte and Pauwels, 2021; Davis et al., 2020), has led to a widespread assumption in the field: that non-differentiability in neural networks is inconsequential (see Section 6.1.5 of Stevens et al.
+(2020), for instance).
+Notably, three of the most widely cited surveys on optimization methods for neural networks (Bottou et al., 2018; Le et al., 2011; Ruder, 2016) do not discuss non-differentiability, despite the ReLU—arguably the most ubiquitous activation function in neural networks—being non-differentiable.
+This assumption has not only rekindled interest in gradient descent dynamics within complex loss landscapes (Ahn et al., 2022b; Dauphin et al., 2014; Reddi et al., 2019), but has also led to a trend where theoretical papers develop results for continuously differentiable functions and then claim universality by running simulations on non-differentiable neural networks (Experiment 6 in Ahn et al.
+(2022b), Section 3.1 in Ma et al.
+(2022), Section 6 in Zhang et al.
+(2022)).
+Our paper examines this fundamental assumption, demonstrating that it leads to significant inaccuracies in several highly referenced papers and texts regarding neural network training dynamics.
+We analyze three critical aspects of the training process, making the following key contributions:
+  • Convergence Analysis of ReLU Networks: We show that NGDMs converge more slowly than GDs, suggesting that many convergence rate claims made under the assumption of $L-smoothness$ have limited applicability to NGDM properties.
+  • Solutions to the LASSO problem: We demonstrate that NGDMs do not produce sparse solutions for the LASSO problem, even in the simplest case of the $L_{1}$ penalized linear model.
+This finding contrasts with prevalent understanding.
+Moreover, we show that NGDMs can yield unexpected results, such as producing solutions with larger $L_{1}$ norms when using larger LASSO penalties.
+Importantly, we prove that variants like RMSProp and NDGM with momentum behave differently from vanilla NDGM, challenging the common assumption that these algorithms converge to similar solutions in non-differentiable settings.
   • The Edge of Stability: We demonstrate that the edge of stability conjecture does not hold, even for gradient descents on all convex differentiable neural networks.
-Furthermore, we provide a counter-example to the claim made in (Ma et al., 2022) that subquadratic behavior around the local minimum is responsible for the Edge of Stability phenomenon.
-In a broader context, our analysis emphasizes the critical role the assumptions we make play on our understanding of these complex systems.
-For NGDMs in particular, our results suggest a need to treat GDs and NGDMs as distinct systems with different dynamics, challenging the common assumption of their similarity.
+Furthermore, we provide a counter-example to the claim made in Ma et al.
+(2022) that subquadratic behavior around the local minimum is responsible for the Edge of Stability phenomenon.
+In a broader context, our analysis underscores the importance of the assumptions we make in understanding these complex systems.
+For NGDMs in particular, our results suggest a need to reconsider how we approach these methods, challenging the common assumption of their similarity to GDs and calling for a more nuanced theoretical foundation in deep learning optimization.
 2 Preliminaries
-In the three sections that follow, we show how disregarding the non-differentiability of the loss function severely limits our understanding of different aspects of training dynamics.
-Each of these sections is largely self-contained, complete with a literature review, problem setup, and analysis.
-Throughout this paper, experiments involving deep networks consistently employ convolutional neural network (CNN) architectures akin to VGG16 (Simonyan & Zisserman, 2014).
-This choice is motivated by the architecture’s widespread usage in the papers referenced throughout our discussions.
-Our experiments have relatively low computational costs, with the most expensive run completing within 2 hours on a single GPU.
-The present section addresses background and terminology that is pertinent to all the subsequent sections.
-NDGM v/s GD:
-The crux of our analyses lies in the structural differences between GDs and NDGMs.
+In the following three sections, we demonstrate how disregarding the non-differentiability of the loss function significantly limits our understanding of various aspects of training dynamics.
+Each section is largely self-contained, including its own literature review, problem setup, and analysis.
+Throughout this paper, our experiments with deep networks consistently employ convolutional neural network (CNN) architectures similar to VGG16 (Simonyan and Zisserman, 2014), chosen for its widespread use in the papers we reference.
+Our experiments have relatively low computational demands, with the most expensive run completing within 2 hours on a single GPU.
+This section introduces background and terminology relevant to all subsequent sections.
+NDGM vs. GD:
+The core of our analyses lies in the structural differences between GDs and NDGMs.
 For a continuously differentiable loss function, $f(\beta)$, the well-studied GD sequence (Bertsekas, 1997; Boyd et al., 2004) is described by the recursion
@@
 where ~$\widetilde{\nabla}g(\tilde{\beta}_{t})=\nabla g(\tilde{\beta}_{t})$ if the gradient exists at $\tilde{\beta}_{t}$, else it is a heuristic measure  [^1] https://rb.gy/g74av.
-The deep learning literature widely assumes that the dynamics described by the recursions in (1) and (2) are nearly the same.
+The deep learning literature widely assumes that the dynamics described by the recursions in (1) and (2) are nearly identical.
 Convex non-differentiable loss functions:
-In this paper, we show the problems with this assumption using a variety of convex loss functions.
-The majority of our analysis is performed on the penalized single layer neural network regression problem with a non-positive response vector, which (Kumar, 2023) has shown to be convex with with the all-zeros vector as its unique global minimizer.
-The loss function for this problem is given by
-$\begin{split}L_{1}(\beta;{\bf{Z}},{\bf{y}},\lambda_{1},\lambda_{2})&=||{\bf{y}}-\max(0,{\bf{Z}}\beta)||^{2}\\ &=+\lambda_{1}||\beta||_{1}+\lambda_{2}||\beta||_{2}^{2},\end{split}$ (3)
-where $\beta$ is a $P\times 1$ parameter vector, ${\bf{Z}}$ is an $N\times P$ data matrix whose $i^{th}$ column is given by $z_{i}$, ${\bf{y}}$ is a $N\times 1$ non-positive response vector, $\lambda_{1}\geq 0$ is the LASSO penalty, $\lambda_{2}\geq 0$ is the ridge penalty, and $\lambda_{1}+\lambda_{2}>0$.
-Throughout this paper, we denote the learning rate by $\alpha$, and the $i^{th}$ entry in the parameter vector after $k$ iterations as $\beta_{k}[i]$, with $10\leq i\leq P-$ 1.
+We demonstrate the issues with this assumption using various convex loss functions.
+Most of our analysis focuses on the penalized single-layer neural network regression problem with a non-positive response vector, which Kumar (2023) has shown to be convex with the all-zeros vector as its unique global minimizer.
+The loss function for this problem is:
+$L_{1}($ $\beta;{\bf{Z}},{\bf{y}},\lambda_{1},\lambda_{2})=$
+$\|{\bf{y}}-\max(0,{\bf{Z}}\beta)\|^{2}+\lambda_{1}\|\beta\|_{1}+\lambda_{2}\|\beta\|_{2}^{2},$ (3)
+where $\beta$ is a $P\times 1$ parameter vector, ${\bf{Z}}$ is an $N\times P$ data matrix with $i^{th}$ column $z_{i}$, ${\bf{y}}$ is a $N\times 1$ non-positive response vector, $\lambda_{1}\geq 0$ is the LASSO penalty, $\lambda_{2}\geq 0$ is the ridge penalty, and $\lambda_{1}+\lambda_{2}>0$.
+We denote the learning rate by $\alpha$, and the $i^{th}$ entry in the parameter vector after $k$ iterations as $\beta_{k}[i]$, with $10\leq i\leq P-$ 1.
 The subgradient method:
-Our analysis heavily relies on the observation that for a convex non-differentiable loss functions, the NDGM recursion described in (2) effectively implements the subgradient method (Bertsekas, 1997; Shor, 2012).
-The theoretical properties of this method are well studied in the optimization literature, with a concise overview of key findings available in the lecture notes of (Boyd et al., 2003) and (Tibshirani, 2015).
-The subgradient method and GD recursions have different properties, and these differences play a key role in our subsequent discussions.
-We reference the relevant portions from the lecture notes as they arise.
+Our analysis relies heavily on the observation that for convex non-differentiable loss functions, the NDGM recursion in (2) effectively implements the subgradient method (Bertsekas, 1997; Shor, 2012).
+The theoretical properties of this method are well-studied in optimization literature, with key findings summarized in the lecture notes of Boyd et al.
+(2003) and Tibshirani (2015).
+The subgradient method and GD recursions have different properties, which play a crucial role in our discussions.
+We will reference relevant portions of these lecture notes as they become pertinent to our analysis.
 Constant, diminishing, and reducing step sizes:
-Our formulations in equations (1) and (2) employ a constant learning rate.
-This choice bears some explaining, for which we first recall some definitions from the optimization literature (see Chapter 1 of (Bertsekas, 1997) for details).
-In the GD literature, the training regimen described in equation (1) is called a constant step-size regime.
-In this regime, the GD is initialized using some randomized approach, and iterated with a constant learning rate until the loss stops decreasing.
-Although constant step-size GDs approach the stationary point, they rarely converge to it unless the loss function possesses desirable properties like $L-smoothness$ (Proposition 1.2.3 in (Bertsekas, 1997)).
-To ensure provable convergence to the stationary point in more general settings, studies often adopt the diminishing step-size regime.
-This regime is characterized by the iteration
+Our formulations in (1) and (2) use a constant learning rate.
+We explain this choice by recalling definitions from optimization literature (see Chapter 1 of Bertsekas (1997)).
+The training regimen in (1) is called a constant step-size regime, where GD is initialized randomly and iterated with a constant learning rate until the loss stops decreasing.
+Constant step-size GDs approach the stationary point but rarely converge to it unless the loss function has desirable properties like $L-smoothness$ (Proposition 1.2.3 in Bertsekas (1997)).
+For provable convergence in more general settings, studies often use the diminishing step-size regime, characterized by:
 $\beta_{t+1}=\beta_{t}-\alpha_{t}\nabla f(\beta_{t}),$ (4)
 where $\lim_{t\rightarrow\infty}\alpha_{t}=0$, and $\sum_{t}\alpha_{t}\rightarrow\infty$.
-Convergence results in this regime are derived under ideal conditions, and provide guarantees that the recursion will reach the stationary point at least in the theoretical limit.
-In practice, diminishing step-size GDs are implemented with a stopping criteria which determines when the solution produced by the GD is good enough to end training.
-The combination of the diminishing step-size regime and stopping criteria commonly used in the neural network literature is what we term the “reducing” step-size regime.
-In this regime, training unfolds as follows: commencing with a weight of $\beta_{0}$, GD is run for some $N_{0}$ iterations with a learning rate of $\alpha_{0}$, followed by $N_{1}$ iterations with a reduced learning rate of $\alpha_{1}<\alpha_{0}$, and so on, until the engineer is satisfied with the solution.
-In the reducing step-size regime, training involves chaining together a finite number (say $T$) of constant step-size GD runs, with the $(i+1)^{th}$ run having an initialization of $\beta_{u}$ and a learning rate of $\alpha_{i}$, where $u=\sum_{j=0}^{i-1}N_{j}$.
-While the theoretical guarantees of the diminishing step-size regime may not be realized, the limiting dynamics of the sequence in this regime are exactly those of a constant step-size GD, with learning rates and weight initializations chosen corresponding to the values at the beginning of the $T^{th}$ run.
-Accordingly, we exclusively study the constant step-size problem, assuming appropriate weights and learning rates have been chosen.
+Convergence results in this regime provide theoretical guarantees under ideal conditions.
+In practice, diminishing step-size GDs use stopping criteria to end training.
+We term the combination of diminishing step-size and stopping criteria commonly used in neural network literature the "reducing" step-size regime.
+In this regime, training proceeds as follows: starting with weight $\beta_{0}$, GD runs for $N_{0}$ iterations with learning rate $\alpha_{0}$, followed by $N_{1}$ iterations with reduced rate $\alpha_{1}<\alpha_{0}$, and so on, until satisfactory results are achieved.
+This involves chaining together $T$ constant step-size GD runs, with the $(i+1)^{th}$ run initialized at $\beta_{u}$ with learning rate $\alpha_{i}$, where $u=\sum_{j=0}^{i-1}N_{j}$.
+While theoretical guarantees of the diminishing step-size regime may not be realized, the limiting dynamics in this regime are those of a constant step-size GD, with learning rates and weight initializations chosen from the $T^{th}$ run.
+Therefore, we exclusively study the constant step-size problem, assuming appropriate weight initializations and learning rates.
 3 NDGM and convergence analysis
-Recent theoretical work by (Davis et al., 2020) has shown that in the diminishing step-size regime, NDGMs converge to a local optimal solution in the theoretical limit under fairly general conditions.
-Unfortunately, the conditions of the diminishing step size regime are rarely satisfied in practice.
-To illustrate this point, consider the Assumption A.3 in (Davis et al., 2020), which requires that in the diminishing step size regime, the sum of the learning rates should diverge.
-As was previously stated, this is a standard assumption even in the thoeretical convergence analysis of GDs.
-However, the prevelant step-size reduction we observed in most training scripts involves decaying the learning rate by a constant multiplicative factor at predefined intervals (StepLR in PyTorch [^2] https://rb.gy/pub4s).
-With this decay scheme, the learning rates form a convergent geometric series, thus violating the assumption.
-For the differentiable case, we know that even when the learning rates form a convergent geometric series, the GD sequence will come within some vicinity of the optimal solution.
-Therefore, the question arises if the same holds true for NDGM sequences.
-More precisely, can we state that the dynamics of NDGM and GD sequences are identical?
+Recent theoretical work by Davis et al.
+(2020) has shown that in the diminishing step-size regime, NDGMs converge to a local optimal solution in the theoretical limit under fairly general conditions.
+However, these conditions are rarely satisfied in practice.
+For instance, Assumption A.3 in Davis et al.
+(2020) requires that the sum of the learning rates should diverge in the diminishing step size regime.
+While this is a standard assumption for GD analysis, it’s often violated in practice.
+The prevalent step-size reduction in most training scripts involves decaying the learning rate by a constant multiplicative factor at predefined intervals (e.g., StepLR in PyTorch [^2] https://rb.gy/pub4s), resulting in a convergent geometric series of learning rates, directly contradicting this assumption.
+For differentiable functions, we know that even with a convergent geometric series of learning rates, the GD sequence approaches the vicinity of the optimal solution.
+However, it’s unclear if this holds true for NDGM sequences.
+More precisely, can we assert that the dynamics of NDGM and GD sequences are identical?
 NDGM and GD dynamics are not identical:
-The dynamics of the two sequences are not identical, and we demonstrate this by arriving at a contradiction.
+We demonstrate the dynamics of NDGM and GD sequences are different by arriving at a contradiction.
 Suppose the sequence $\{\beta_{k}\}_{k=0}^{\infty}$ obtained by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})=L_{1}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1},0,0.01)$ has the same properties as a GD sequence.
-Then, the entire NDGM sequence can be characterized using the ”Capture Theorem” as described below
+Then, the entire NDGM sequence can be characterized using the "Capture Theorem":
 Proposition 3.1.
-Let $x_{k+1}=x_{k}-\alpha\nabla f(x_{k})$ be a sequence generated by running GD on a continuously differentiable function, $f(x)$ such that $f(x_{k+1})\leq f(x_{k})$.
+Let $x_{k+1}=x_{k}-\alpha\nabla f(x_{k})$ be a sequence generated by running GD on a continuously differentiable function, $f(x)$ such that $f(x_{k+1})\leq f(x_{k})$ for all $k$.
 If $x^{*}$ is an isolated local minimum of $f(x)$ in a unit ball around $x^{*}$, and $||x_{0}-x^{*}||<1$, then $||x_{k}-x^{*}||<1$ for all $k>0$.
 Proof.
-Result follows by using $\bar{\epsilon}=1$ in Proposition 1.2.5 of (Bertsekas, 1997). ∎
-Therefore, if $||\beta_{0}||_{\infty}<1$, and $L_{2}(\beta_{k+1};{\bf{Z}}_{1},{\bf{q}}_{1})\leq L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$ for all $k\geq 0$, then we expect $||\beta_{k}||_{\infty}<1$ for all $k\geq 0$.
-To validate this claim, we first generate a $20\times 500$ matrix ${\bf{Z}}_{1}$ with entries uniformly sampled on $[-1,1]$.
-Next, we generate a $20\times 1$ vector ${\bf{q}_{1}}$ with entries uniformly sampled on $[-1,0]$.
-Finally, we generate the sequence $\{\beta_{t}\}_{t=0}^{\infty}$, by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})$ with the entries in $\beta_{0}$ uniformly sampled on $[-1,1]$.
-As seen in Figure 1, although the loss function is a monotone decreasing function of the iteration number, the infinity norm of the NDGM sequence eventually takes values greater than 1, hence establishing the contradiction.
-The non-capturing described in Figure 1 is to be expected; as stated in Section 1 of (Boyd et al., 2003), ”Unlike the ordinary gradient method, the subgradient method is not a descent method; the function value can (and often does) increase”.
-Figure 1: Plot of the loss function ($L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$), and infinity norm ($||\beta_{k}||_{\infty}$) of the parameter vector as a function of the iteration number ($k$).
-Although the loss function is monotone decreasing, the infinity norm of the weights does not ”capture”, hence demonstrating that the dynamics of NDGM and GD are not identical
+Result follows by using $\bar{\epsilon}=1$ in Proposition 1.2.5 of Bertsekas (1997). ∎
+Therefore, if $||\beta_{0}||_{\infty}<1$, and $L_{2}(\beta_{k+1};{\bf{Z}}_{1},{\bf{q}}_{1})\leq L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$ for all $k\geq 0$, we expect $||\beta_{k}||_{\infty}<1$ for all $k\geq 0$.
+To validate this claim, we generate a $20\times 500$ matrix ${\bf{Z}}_{1}$ with entries uniformly sampled on $[-1,1]$, a $20\times 1$ vector ${\bf{q}_{1}}$ with entries uniformly sampled on $[-1,0]$, and the sequence $\{\beta_{t}\}_{t=0}^{\infty}$ by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})$ with entries in $\beta_{0}$ uniformly sampled on $[-1,1]$.
+As shown in Figure 1, although the loss function decreases monotonically with iteration number, the infinity norm of the NDGM sequence eventually exceeds 1, contradicting our assumption.
+This non-capturing behavior is expected; as Boyd et al.
+(2003) note, "Unlike the ordinary gradient method, the subgradient method is not a descent method; the function value can (and often does) increase".
+Refer to captionRefer to caption
+Figure 1: Loss Function and Parameter Norm During Training.
+Left: Loss function $L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$ vs.
+iteration $k$.
+Right: Infinity norm $\|\beta_{k}\|_{\infty}$ vs.
+iteration $k$.
+The monotonically decreasing loss contrasts with the non-monotonic parameter norm, indicating differing dynamics between NDGM and GD.
 Convergence analysis and L-smoothness:
-Given that the dynamics of GDs and NDGMs differ, the question arises what assumptions are appropriate for the study of convergence analysis in neural network training.
-Majority of the analyses in the literature ((Bottou et al., 2018; Zhang et al., 2022; Défossez et al., 2020; You et al., 2019) to state a few) assume that the loss function, $f(\beta)$ is continuously differentiable, and is $L-smooth$, i.e., there exists a constant $L$, such that
+Given the distinct dynamics of GDs and NDGMs, we must reconsider the assumptions appropriate for convergence analysis in neural network training.
+Many analyses in the literature (Bottou et al., 2018; Zhang et al., 2022; Défossez et al., 2020; You et al., 2019) assume that the loss function $f(\beta)$ is continuously differentiable and $L-smooth$, i.e., there exists a constant $L$ such that:
 $||\nabla f(x)-\nabla f(y)||\leq L||x-y||$ (5)
 for all $x$, $y$ in the domain of $f$.
-However, this assumption breaks down for non-differentiable functions, where the gradient does not even exist at various points in the domain.
-This raises the question if the assumption can be salvaged by confining its validity to a local context around the local optimal solution.
-Unfortunately, even this cannot be done; as stated in (Laurent & Brecht, 2018), “local minima of ReLU networks are generically nondifferentiable.
-They cannot be waved away as a technicality, so any study of the loss surface of such network must invoke nonsmooth analysis”.
-Moreover, assuming that the loss function is $L-smooth$ grossly over-estimates the speed of convergence of NDGMs.
-To illustrate this, consider that in order to achieve an error level of $\epsilon$, GD applied to an $L-smooth$ convex loss function, $f$, would require $O(1/\epsilon)$ steps for convergence.
-In contrast, NDGM applied to a convex non-differentiable loss function, $g$, satisfying the Lipschitz condition
+However, this assumption fails for non-differentiable functions where the gradient doesn’t exist at various points.
+One might consider salvaging this assumption by limiting its validity to a local context around the local optimal solution.
+Unfortunately, even this approach is untenable.
+As Laurent and Brecht (2018) state, "local minima of ReLU networks are generically nondifferentiable.
+They cannot be waved away as a technicality, so any study of the loss surface of such network must invoke nonsmooth analysis".
+This observation aligns with established ideas in convex optimization.
+For instance, the development and application of second-order cone programs (SOCPs) for convex non-smooth problems was motivated by similar considerations (Boyd et al., 2004) [^3] See 26:30 to 29:00 in https://rb.gy/ab0aeh.
+In both SOCPs and ReLU networks, local minima often occur at points of non-differentiability, necessitating non-smooth analysis.
+Moreover, assuming $L-smoothness$ significantly overestimates the convergence speed of NDGMs.
+For an error level of $\epsilon$, GD applied to an $L-smooth$ convex loss function requires $O(1/\epsilon)$ steps for convergence.
+In contrast, NDGM applied to a convex non-differentiable loss function satisfying the Lipschitz condition:
 $||g(x)-g(y)||\leq L||x-y||$ (6)
-would take $O(1/\epsilon^{2})$ steps (see Section 7.1.3 of (Tibshirani, 2015) for the result, and a visual illustration of how much slower $O(1/\epsilon^{2})$ is than $O(1/\epsilon)$).
-Thus, it is likely that the vast literature on the convergence analysis of GDs for loss functions satisfying (5) have little bearing on the dynamics of convergence in modern non-smooth neural networks.
-More generally, we expect the speed of convergence for smooth neural networks to be faster than that of non-smooth neural networks.
+takes $O(1/\epsilon^{2})$ steps (see Section 7.1.3 of Tibshirani (2015) for the result and a visual illustration of how much slower $O(1/\epsilon^{2})$ is than $O(1/\epsilon)$).
+Thus, the extensive literature on GD convergence analysis for $L-smooth$ functions likely has limited relevance to the convergence dynamics of modern non-smooth neural networks.
+Generally, we expect smooth neural networks to converge faster than non-smooth ones.
 Experiments with NDGM convergence:
-We conduct two sets of experiments on CNNs of varying depths to validate our intuition in a more general context (see the Appendix A for architecture details).
-In each set, we perform two identical simulations for every network, with the only difference being the choice of activation function: one simulation uses the the non-smooth ReLU activation function, while the other uses the smooth and asymptotically equivalent GELU activation function (Hendrycks & Gimpel, 2016).
+We conduct two sets of experiments on CNNs of varying depths to validate our intuition in a more general context (see the Appendix A.1 for architecture details).
+In each set, we perform two identical simulations for every network, with the only difference being the choice of activation function: one simulation uses the non-smooth ReLU activation function, while the other uses the smooth and asymptotically equivalent GELU activation function (Hendrycks and Gimpel, 2016).
 The first set of experiments excludes batch normalization, while the second includes it.
 Figure 2 shows that without batch normalization, the gap in convergence speeds between smooth and non-smooth networks widens as network depth increases, aligning with our expectations.
-This trend is further substantiated by (Hendrycks & Gimpel, 2016), who demonstrate accelerated convergence in GELU networks compared to ReLU networks across diverse architectures and datasets (see Figures 2 through 7 in their paper).
+This trend is further substantiated by Hendrycks and Gimpel (2016), who demonstrate accelerated convergence in GELU networks compared to ReLU networks across diverse architectures and datasets (see Figures 2 through 7 in their paper).
 However, with batch normalization, the difference between ReLU and GELU training curves becomes indistinguishable, emphasizing its importance in the training process.
 We conclude that the choice of activation function is the primary factor contributing to the difference in convergence speeds between ReLU and GELU curves with increasing depth, as no other principled reason appears evident.
-Figure 2: Training curves comparing smooth and non-smooth CNNs (from left to right) of depth 11, 13, and 19, respectively.
-The top row shows validation losses without batch normalization, while the bottom row includes batch normalization.
+Refer to captionRefer to captionRefer to captionRefer to caption
+Figure 2: Training curves comparing smooth and non-smooth CNNs of depth 11 (left), and 19 (right). Top: without batch normalization. Bottom: with batch normalization.
 L-smoothness and the geometry of ReLU networks:
-As stated in (Bottou, 2010), the assumption of $L-smoothness$ ensures that the gradient of the loss function does not change too rapidly with respect to the parameter vector.
-However, this assumption which does not hold for ReLU networks.
+The widely used $L-smoothness$ assumption, which ensures that the gradient of the loss function does not change too rapidly with respect to the parameter vector (Section 4.1 in Bottou et al.
+(2018)), does not hold for ReLU networks.
 To illustrate this, consider the 1-D case with $q(x)=\text{max}(0,x)$.
 In this case, $\lim_{h\rightarrow 0^{+}}(1/2h)(\nabla q(h)-\nabla q(-h))=\infty$, indicating that the gradient changes arbitrarily quickly at the origin.
-This observation necessitates the use of non-smooth analysis for studying the loss functions of ReLU networks, as detailed in (Laurent & Brecht, 2018).
-The key intuition about non-differentiable functions is that they have kinks corresponding to non-differentiable points, and local minima frequently lie at these kinks.
-As a result, gradient descents which rely on smoothness assumptions cannot access these points of interest.
-This realization serves as one of the primary motivations for the development and application of second-order cone programs for convex non-smooth problems (Boyd et al., 2004)  [^3] See 26:30 to 29:00 in https://rb.gy/ab0aeh.
-The inability of gradient descent methods to effectively maneuver non-smooth functions and the associated pathologies form a recurring theme throughout the remaining sections of this paper.
-In a subsequent section on the Edge of Stability, we show that insights based on the assumption of $L-smoothness$ are not universally applicable even across all Lipschitz continuous, convex, differentiable neural networks.
+This necessitates the use of non-smooth analysis for studying these losses.
+The issue, as previously discussed, is that non-differentiable functions have kinks corresponding to non-differentiable points, and local minima frequently lie at these kinks.
+Therefore, gradient-based approaches developed for differentiable functions are insufficient since they cannot access these points of interest (see Appendix A.2 for geometric intuitions into this fact).
+The inability of gradient descent methods to effectively maneuver non-smooth functions and the associated pathologies form a recurring theme in this paper.
 $L-smoothness$ is one of the strongest assumptions after convexity, guaranteeing the existence and boundedness of the first and second derivatives—criteria unmet by most loss functions in neural network literature.
+To emphasize this point, we show in a later section that results derived for smooth networks may not generalize even to all Lipschitz continuous convex differentiable networks.
 Limitations of our work:
 A common limitation in the current and subsequent sections is that we do not analyze the dynamics of the neural network training process as it is often conducted in practice.
-This is necessitated by the fact that neural networks often involve non-differentiable layers like dropout and max pooling, which are beyond the scope of our study.
+This is necessitated by the fact that neural networks often involve dropout layers, which leave the loss function discontinuous, let alone non-differentiable, and are thus beyond the scope of our study.
 We do not reiterate this limitation in every subsequent section.
+Extending our results to the general neural network setting:
+Our analysis in 3 focuses on the specific case of non-positive response vectors, which allows for a convex formulation.
+Extending these results to general single-layer neural networks with arbitrary response vectors, let alone multi-layer networks, is challenging due to the resulting non-convex and non-differentiable loss functions.
+Currently, there is no known way to rigorously analyze such networks in their full generality.
+This limitation is not unique to our work—it is a fundamental challenge in the field.
+Our paper’s key contribution is to rigorously demonstrate that this challenge cannot be addressed by simply extrapolating results from the analysis of smooth functions to non-smooth settings, as is often done in the literature.
+We specifically caution against the common practice of deriving results for $L-smooth$ networks and then claiming generality based on simulations on architectures like VGG16 or ResNet.
+This approach, while prevalent, fails to capture the fundamental differences between smooth and non-smooth optimization that we highlight.
+Our findings are crucial for the field, as they expose the limitations of current methods and underscore the need for more rigorous and appropriate techniques in analyzing ReLU networks and other non-smooth architectures.
 The way forward:
 In this section, we show that NDGMs are slower to converge than GDs.
-Fortunately, most recent explorations into image and language models primarily leverage the transformer architecture (Vaswani et al., 2017) with the differentiable GELU activation function (Hendrycks & Gimpel, 2016).
+Fortunately, most recent explorations into image and language models primarily leverage the transformer architecture (Vaswani et al., 2017) with the differentiable GELU activation function (Hendrycks and Gimpel, 2016).
 As a consequence, future iterations of neural networks may predominantly employ pure GDs.
@@
 Zero weights are crucial for model compression (Blalock et al., 2020; Han et al., 2015a, b; Li et al., 2016), and inducing sparsity by penalizing the $L_{1}$ norm of model weights is a common approach.
-While the statistics literature employs specialized algorithms for problems involving penalized $L_{1}$ norms (Efron et al., 2004; Friedman et al., 2008; Mazumder & Hastie, 2012; Tibshirani, 1996; Tibshirani et al., 2005; Zou & Hastie, 2005), neural networks often use the penalized $L_{1}$ norm with arbitrary non-convex functions, assuming that NDGM results in a near-sparse optimal solution (Bengio, 2012; Goodfellow et al., 2016b; Scardapane et al., 2017).
-This expectation is not always met, as evidenced by user struggles in PyTorch forums [^4] https://rb.gy/kh0qr [^5] https://rb.gy/2vsry [^6] https://rb.gy/6dfli.
-Given the LASSO’s critical applications (Ghosh et al., 2021; Li et al., 2011; Li & Sillanpää, 2012; Ogutu et al., 2012), understanding the reliability of the NDGM solution for the LASSO problem is essential.
+While the statistics literature employs specialized algorithms for problems involving penalized $L_{1}$ norms (Efron et al., 2004; Friedman et al., 2008; Mazumder and Hastie, 2012; Tibshirani, 1996; Tibshirani et al., 2005; Zou and Hastie, 2005), neural networks often use the penalized $L_{1}$ norm with arbitrary non-convex functions, assuming that NDGM results in a near-sparse optimal solution (Bengio, 2012; Goodfellow et al., 2016b; Scardapane et al., 2017).
+However, this assumption is not satisfied in practice, as evidenced by user struggles in PyTorch forums [^4] https://rb.gy/kh0qr [^5] https://rb.gy/2vsry [^6] https://rb.gy/6dfli.
+Given the LASSO’s critical applications (Ghosh et al., 2021; Li et al., 2011; Li and Sillanpää, 2012; Ogutu et al., 2012), understanding the reliability of the NDGM solution for the LASSO problem is essential.
 NDGM solutions for the LASSO problem are not reliable:
-We show that NDGM produces unreliable solutions even for the simple case of the LASSO penalized linear model.
-In particular, we show that increasing the LASSO penalty can result in a NDGM solution with a larger $L_{1}$ norm, completely defeating the purpose of the penalty.
+We demonstrate that NDGM produces unreliable solutions even for the simple case of the LASSO penalized linear model.
+Specifically, we show that increasing the LASSO penalty can result in a NDGM solution with a larger $L_{1}$ norm, completely defeating the purpose of the penalty.
 We first analytically demonstrate the unreliability for the simplest LASSO problem,
@@
 Furthermore, there exists an integer $N_{0}$, such that for every $N\geq N_{0}$, we can write
-$\beta_{n_{k}+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even},\end{cases}$ (8)
+$\beta_{N+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even},\end{cases}$ (8)
 where $\gamma_{k}\in(0,\alpha\lambda_{1})$ for all $10\leq k\leq P-$ 1.
 Proof.
-See Appendix B for details ∎
+See Appendix B.1 for details ∎
 Equation (8) suggests that with the same learning rate and initialization, a larger $\lambda_{1}$ can lead to a NDGM solution with a larger $L_{1}$ norm, defeating the purpose of the LASSO penalty.
 To illustrate this, we note that in the 2D case, with weights initialized as $\beta_{0}=[0.5053,0.5053]$, $\alpha=0.01$, and $\lambda_{1}=1$, the NDGM sequence oscillates between $[0.0053,0.0053]$ and $[-0.0047,-0.0047]$.
-However, with $\lambda_{1}=100$, and the same values of $\beta_{0}$ and $\alpha_{1}$, the solution oscillates between $[0.5053,0.5053]$ and $[-0.4947,-0.4947]$.
+However, with $\lambda_{1}=100$, and the same values of $\beta_{0}$ and $\alpha$, the solution oscillates between $[0.5053,0.5053]$ and $[-0.4947,-0.4947]$.
 Therefore at convergence, the $L_{1}$ norm is at most 0.0106 for $\lambda_{1}=1$, but at least 0.9894 for $\lambda_{1}=100$.
@@
 With $\lambda_{1}=1$, the sequence oscillates between $[0.0003,0.0003]$ and $[-0.0007,-0.0007]$, while with $\lambda_{1}=100$, it oscillates between $[0.0053,0.0053]$ and $[-0.0947,-0.0947]$.
-Appendix C shows that this paradox holds for NDGM solutions of the standard $L_{1}$ regularized linear problem.
+Appendix B.2 shows via simulation that this paradox holds for NDGM solutions of the standard $L_{1}$ regularized linear problem.
 The subgradient method and step-size choices:
 The counter-intuitive behavior seen in the example above can be explained using the subgradient method.
-Since the loss function described in equation (7) is Lipschitz continuous (equation (6)) with $L=\lambda_{1}$, we have from (Boyd et al., 2003) (Section 2 on the Constant Step Size) that
+Since the loss function described in equation (7) is Lipschitz continuous (equation (6)) with $L=\lambda_{1}$, we have from Boyd et al.
+(2003) (Section 2 on the Constant Step Size) that
 $\lim_{k\rightarrow\infty}||\beta_{k}||_{1}<\alpha\lambda_{1}.$ (9)
@@
 Connections to other NDGM variants:
-Conventional wisdom in deep learning theory suggests that different NDGM variants, such as NDGM with momentum and RMSProp, converge at varying rates but eventually reach similar near-optimal values (Figure 8.5 of (Goodfellow et al., 2016a), slides 64 onwards in https://rb.gy/ojzpgm  [^7] These slides are from one of the most widely used courses for self teaching; see 1:01:00 - 1:03:00 of https://rb.gy/kaj976 where this precise point is made).
-However, we prove in Appendices D and E that these claims are incorrect even for the toy LASSO problem described in (7).
-Furthermore, our analysis reveals two interesting facts.
-First, that the RMSProp sequence for (7) is agnostic of the value of $\lambda_{1}$ used in the analysis.
-Second, there is a phase transition between the RMSProp and vanilla NDGM training curves characterized by a change in the relative behavior of the training curves of the two algorithms as $\lambda_{1}$ is increased from a value which is much smaller than 1 to a value much larger than 1, which cannot be explained without accounting for the non-differentiability of the $L_{1}$ regularizer.
+Conventional wisdom in deep learning theory suggests that different NDGM variants converge at varying rates but eventually reach similar near-optimal values (Figure 8.5 of Goodfellow et al.
+(2016a), slides 64 onwards in https://rb.gy/ojzpgm  [^7] These slides are from one of the most widely used courses for self teaching; see 1:01:00 - 1:03:00 of https://rb.gy/kaj976 where this precise point is made).
+However, we analytically show that these claims are incorrect even for the toy LASSO problem described in (7) (see Appendices B.3 and B.4 for details).
+Furthermore, our analysis reveals two interesting facts which cannot be explained by the analysis of differentiable functions: 1) the RMSProp sequence for (7) is agnostic of the value of $\lambda_{1}$ used in the analysis, and 2) there is a phase transition between the RMSProp and NDGM training curves based on the magnitude of $\lambda_{1}$.
 These findings highlight the importance of incorporating the non-differentiability of regularizers like the $L_{1}$ norm when analyzing the behavior of NDGM variants.
-Failing to account for non-differentiability results in an incomplete understanding of the optimization dynamics.
-Consequently, it remains unclear how the convergence analysis of algorithms like Adam and RMSProp for differentiable loss functions, as often performed in studies such as (Reddi et al., 2019; Défossez et al., 2020; Zhang et al., 2022), relates to the dynamics of these algorithms’ convergence in non-differentiable settings involving regularizers like the $L_{1}$ norm or activation functions like the ReLU.
 Experiments with deep networks:
-In their seminal work on network pruning, (Han et al., 2015b) propose using $L_{1}$ regularization to encourage weights to approach 0, facilitating the pruning process.
-This approach is widely adopted in the literature (Wen et al., 2016; Xiao et al., 2019; Yang et al., 2019) and industry [^8] See the recommendation of the regularizer in the Training config section of https://rb.gy/20y13m and https://rb.gy/i1sb6s.
-To challenge this suggestion’s general applicability, we conduct two sets of experiments finetuning a pre-trained VGG16 model on CIFAR-10 using $L_{1}$ penalized cross-entropy loss with a constant learning rate of $\alpha=0.003$ for 200 epochs.
-We use $\lambda_{1}=0.0002$ in the first set and $\lambda_{1}=2$ in the second, running three parallel trials within each set using RMSProp, SGD, and Adam.
-At the end of our runs, we make two key observations (Figure 5): 1) The $L_{1}$ norms at convergence differ vastly across the three variants, and 2) for SGD and Adam, larger values of $\lambda_{1}$ result in a larger $L_{1}$ norm at convergence.
-These findings suggest that the relationship between $L_{1}$ regularization and prunable parameters may not be as straightforward as previously thought.
-Indeed, our results show that a larger $\lambda_{1}$ does not necessarily correspond to more prunable parameters (Table 1): with $\lambda_{1}=2$, less than 0.2% of the parameters at convergence with vanilla NDGM have absolute value less than $10^{-5}$, but with $\lambda_{1}=0.002$, 99.98% do.
+The seminal work on network pruning by Han et al.
+(2015b) proposes using $L_{1}$ regularization to encourage weights to approach 0, facilitating the pruning process.
+This approach has been widely adopted in the literature (Wen et al., 2016; Xiao et al., 2019; Yang et al., 2019) and industry [^8] See the recommendation of the regularizer in the Training config section of https://rb.gy/y7o7hw.
+To examine the general applicability of this method, we conducted two sets of experiments finetuning a pre-trained VGG16 model on CIFAR-10 using $L_{1}$ penalized cross-entropy loss with a constant learning rate of $\alpha=0.003$ for 200 epochs.
+We used $\lambda_{1}=0.0002$ in the first set and $\lambda_{1}=2$ in the second, running three parallel trials within each set using RMSProp, SGD, and Adam.
+Our experiments yielded two key observations (Figure 3): 1) The $L_{1}$ norms at convergence differ substantially across the three optimizer variants, and 2) for SGD and Adam, larger values of $\lambda_{1}$ result in a larger $L_{1}$ norm at convergence.
+These findings suggest that the relationship between $L_{1}$ regularization and prunable parameters may be more complex than previously understood.
+Indeed, our results indicate that a larger $\lambda_{1}$ does not necessarily correspond to more prunable parameters (Table 2): with $\lambda_{1}=2$, less than 0.2% of the parameters at convergence with vanilla NDGM have absolute value less than $10^{-5}$, but with $\lambda_{1}=0.002$, 99.98% do.
+These observations raise important questions about the efficacy and generalizability of $L_{1}$ regularization for network pruning across different optimization algorithms and regularization strengths.
+Refer to captionRefer to caption
+Figure 3: $L_{1}$ norm of the weights as a function of iteration number for $\lambda_{1}=0.002$ (left) and $\lambda_{1}=2$ (right).
+In each figure, the blue, red, and green lines represent curves for RMSProp, Adam, and vanilla SGD respectively.
 Soft-thresholding and ISTA:
 The observation that higher $\lambda_{1}$ values do not yield more prunable parameters indicates that while soft thresholding produces sparse solutions for NDGMs, it does not find optimal solutions.
-This issue extends beyond NDGM to general neural network problems, where any soft thresholding solution yields sparsity but not optimality.
-The root cause is that soft thresholding implicitly solves a generalized ISTA problem (Chen et al., 2023), which assumes the loss function is $L-smooth$ (point 1, Section 3.2 of (Chen et al., 2023)) - an assumption that fails to capture the non-convex and non-differentiable dynamics in general neural network settings, as discussed previously.
+This issue extends to general neural network problems, where any soft thresholding solution yields sparsity but not optimality.
+The root cause is that soft thresholding implicitly solves a generalized ISTA problem (Chen et al., 2023), which assumes the loss function is $L-smooth$ (point 1, Section 3.2 of Chen et al.
+(2023)) - an assumption that fails to capture the non-convex and non-differentiable dynamics in general neural network settings, as discussed previously.
 Consequently, applying the generalized ISTA formulation relying on $L-smoothness$ to neural networks necessarily leads to suboptimal solutions.
@@
 The realization that NDGMs lack this capability predates the widespread adoption of deep learning (Xiao, 2009).
-Despite this, a misconception persists in the deep learning literature, suggesting that such algorithms yield sparse solutions comparable to those from traditional machine learning frameworks like glmnet (Friedman et al., 2021) or scikit-learn (Kramer & Kramer, 2016).
+Despite this, a misconception persists in the deep learning literature, suggesting that such algorithms yield sparse solutions comparable to those from traditional machine learning frameworks like glmnet (Friedman et al., 2021) or scikit-learn (Kramer, 2016).
 However, there is a fundamental disparity between these frameworks: glmnet focuses on a narrow set of problems, guaranteeing optimality, while PyTorch (Paszke et al., 2017) provides flexibility in model creation, but lacks similar guarantees.
 Attaining a solution at a non-differentiable point is challenging, emphasizing the importance of exploring related studies and tools rather than relying on simplistic implementations when precision is crucial.
+Connections to the Edge of Stability:
 Proposition 4.1 shows that the NDGM sequence for the non-differentiable, Lipschitz continuous, convex loss function described in (7) will not diverge to $\infty$ for any finite value of $\alpha$.
@@
 5 NDGM and the Edge of Stability
-In recent years, there has been a growing interest on the topic of “unstable convergence”.
-The literature on the subject (Ahn et al., 2022b, a; Arora et al., 2022; Chen & Bruna, 2022; Cohen et al., 2021; Li et al., 2022) is motivated by the finding that unlike convex quadratic forms, gradient descent on neural networks do not diverge to $\infty$ even when the learning rate is greater than $\alpha^{*}=2/\eta$, where $\eta$ is the dominant eigenvalue of the loss function’s Hessian.
-Instead, the loss function has been empirically demonstrated to converge “unstably”, reducing non-montonically on the long run.
-$\alpha=\alpha^{*}$ is referred to in the literature as the ”Edge of Stability”, since it separates regions of ”stable convergence” ($\alpha<\alpha^{*}$) from regions of “unstable convergence” [^9] We define unstable convergence to mean “non-divergence to $\infty$”.
-Our definition is adapted from Definition 1.1 of (Arora et al., 2022).
-However, we deviate from their definition due to its dependence on the existence of the Hessian at all points between the current and next iterate, a condition that need not be satisfied in non-smooth neural networks.
-($\alpha>\alpha^{*}$).
-The derivation of the Edge of Stability condition assumes that the loss function, $f$, is $L-smooth$ (see equation (5)).
-It is conjectured that neural networks demonstrate unstable convergence because the phenomenon is inconsistent with widespread presumptions in the field of optimization (Cohen et al., 2021; Ahn et al., 2024).
-This conjecture is incorrect for Lipschitz continuous, differentiable, convex functions satisfying (6) as shown in the proposition below.
+In recent years, there has been growing interest in the phenomenon of "unstable convergence" in neural network optimization.
+The literature on this subject (Ahn et al., 2022b, a; Arora et al., 2022; Chen and Bruna, 2022; Cohen et al., 2021; Li et al., 2022) is motivated by an intriguing observation: unlike convex quadratic forms, gradient descent on neural networks does not diverge to $\infty$ even when the learning rate exceeds $\alpha^{*}=2/\eta$, where $\eta$ is the dominant eigenvalue of the loss function’s Hessian.
+Instead, the loss function has been empirically shown to converge "unstably", reducing non-monotonically in the long run.
+The value $\alpha=\alpha^{*}$ is termed the "Edge of Stability" in the literature, as it demarcates regions of "stable convergence" ($\alpha<\alpha^{*}$) from regions of "unstable convergence" ($\alpha>\alpha^{*}$).
+We define unstable convergence as "non-divergence to $\infty$".
+This definition is adapted from Definition 1.1 and the subsequent discussion in Arora et al.
+(2022), which explains why using sharpness greater than $2/\eta$ is not part of the definition.
+We deviate from their definition due to its reliance on the Hessian existing at all points between the current and next iterate, a condition unmet in non-smooth neural networks.
+Importantly, the derivation of the Edge of Stability condition assumes that the loss function, $f$, is $L-smooth$ (see equation (5)).
+It has been conjectured that neural networks exhibit unstable convergence because this phenomenon lies beyond the scope of classical optimization theory (Cohen et al., 2021; Ahn et al., 2024).
+However, we show that this conjecture is incorrect for a broader class of functions, specifically Lipschitz continuous, non-smooth convex functions satisfying (6), as demonstrated in the following proposition:
 Proposition 5.1.
@@
 Proof.
-Note that any loss function satisfying (6) has subgradients bounded above by $L$.
-Additionally, as outlined in Section 1 of Boyd et al.
-(2003), the deductions derived from the subgradient method hold equivalently for differentiable functions.
-This equivalence arises from the fact that the only admissible value for the subgradient of a differentiable convex function at any given point is, indeed, its gradient.
-Accordingly, using the convergence properties of the subgradient methods with constant step sizes from Section 2 of Boyd et al.
-(2003), we have that $\lim_{k\rightarrow\infty}f(x_{k})-f^{*}\leq\alpha L^{2}<\infty$, hence the result.
-∎
-Figure 3: Training curves for $\alpha=0.1$, $\alpha=1$, and $\alpha=10$ with the Lipschitz continuous Huber loss function.
+See Appendix D.1 for details. ∎
+The toy LASSO problem studied in (7) and the Huber loss function, widely used in robust regression and object detection models (Girshick, 2015; Liu et al., 2016; Ren et al., 2015), exemplify non-smooth convex loss functions satisfying Proposition 5.1.
+The LASSO problem is non-differentiable, while the Huber loss function is differentiable but non-smooth, as it is once but not twice differentiable.
+The Huber loss for regressing an arbitrary $50\times 1$ response vector, on a $50\times 200$ data matrix is given by
+$L_{7}(\beta)$ $=\frac{1}{50}\sum l(i),\quad\text{where}$ (10)
+$l(i)$ =$=\begin{cases}\frac{1}{2}(y_{i}-z_{i}^{T}\beta)^{2}&\text{if }\left|(y_{i}-z_{i}^{T}\beta)\right|<1\\ \left(|y_{i}-z_{i}^{T}\beta|-\frac{1}{2}\right)&\text{otherwise}\end{cases}$
+Figure 4 illustrates that the NDGM sequence for (10) does not diverge toward $\infty$, even with a high learning rate such as $\alpha=10$, aligning with our expectations.
+Refer to captionRefer to caption
+Figure 4: Training curves for $\alpha=0.1$ and $\alpha=10$ with the Lipschitz continuous Huber loss function.
 As the learning rate increases, the loss starts oscillating with larger frequencies, but never diverges to $\infty$.
 This is to be expected from Proposition 5.1, since the Huber loss function is convex, non-smooth and has bounded subgradients.
-The toy LASSO problem studied in (7), for which we analytically demonstrated unstable convergence in the previous section, is an example of a non-differentiable convex loss function satisfying Proposition 4.1.
-An example of a differentiable non-smooth convex function satisfying Proposition 4.1 is the Huber loss function, which finds use in many state of the art object detection models (Girshick, 2015; Liu et al., 2016; Ren et al., 2015).
-The Huber loss function is non-smooth because it is once but not twice differentiable.
-For a regression problem involving an arbitrary $50\times 1$ response vector, $\mathbf{y}=[y_{1},y_{2}\ldots y_{N}]^{T}$, and an arbitrary $50\times 200$ data matrix, $\mathbf{Z}$, the Huber loss function is defined as
-$L_{7}(\beta)$ $=\frac{1}{N}\sum l(i),\quad\text{where}$ (10)
-$l(i)$ =$=\begin{cases}\frac{1}{2}(y_{i}-z_{i}^{T}\beta)^{2}&\text{if }\left|(y_{i}-z_{i}^{T}\beta)\right|<1\\ \left(|y_{i}-z_{i}^{T}\beta|-\frac{1}{2}\right)&\text{otherwise}\end{cases}$
-where $z_{i}$ is the $i^{th}$ column of $\mathbf{Z}$.
-Figure 3 shows that the NDGM sequence for (10) does not diverge toward $\infty$, even with a high learning rate, such as $\alpha=10$, which aligns with our expectations.
-Proposition 4.1 suggest that it is unclear and the examples that follow show that the common approach of extrapolating results developed for $L-smooth$ loss functions to the non-smooth or non-differentiable case is problematic.
-To emphasize this point, we provide an example of such an extrapolation in the literature which breaks down in the non-smooth case.
 Subquadratic losses and unstable convergence:
-In section 3.1 of (Ma et al., 2022), the authors use second-order finite differences to analyze the curvature of VGG16 and ResNet near the local minimum, concluding that subquadratic growth explains the edge of stability phenomenon.
-However, their analysis is flawed because ReLU networks have generically non-differentiable local minima (Laurent & Brecht, 2018), where their approximation breaks down.
-The Huber loss function described in equation (10) challenges their statement’s broad applicability.
-This overparameterized convex function has a global minimum of 0, exhibiting quadratic growth nearby.
-(Ma et al., 2022)’s result suggests stable convergence, but Proposition 4 guarantees unstable convergence (Figure 3).
-The discrepancy arises because the Huber loss function lacks the twice differentiability required by Definition 4.1 in (Ma et al., 2022).
-In contrast, $L-smooth$ functions, which are twice differentiable by definition, exhibit subquadratic growth near their local minimum during unstable convergence.
-This example highlights the challenges of extending results developed under strong assumptions like $L-smoothness$ to the general case.
+In section 3.1 of Ma et al.
+(2022), the authors use second-order finite differences to analyze the curvature of VGG16 and ResNet near local minima, concluding that subquadratic growth explains the edge of stability phenomenon.
+However, their analysis is fundamentally flawed for these ReLU networks because second-order finite differences break down at their generically non-differentiable local minima (Laurent and Brecht, 2018).
+The overparameterized Huber loss function (Equation 10) further challenges the general applicability of their claims: despite having a global minimum of 0 with quadratic growth nearby, it exhibits unstable convergence as guaranteed by Proposition 5.1 (Figure 4).
+This discrepancy arises because the Huber loss lacks the twice differentiability required by Definition 4.1 in Ma et al.
+(2022).
+In contrast, all $L-smooth$ functions, being twice differentiable by definition, will show subquadratic growth near their local minima during unstable convergence, aligning with their results.
+This example underscores the risks of extending results derived for smooth functions to non-smooth cases.
 The way forward:
-We demonstrate that our understanding of the Edge of Stability will be incomplete until we more precisely quantify conditions under which we expect it to occur.
-Furthermore, considering ReLU networks’ non-differentiable minima, without further research into NDGM behavior around non-differentiable points our grasp of training dynamics lacks completeness.
-Consequently, we propose employing the convex non-differentiable single-layer neural network presented in (Kumar, 2023) as the toy problem for theory development.
-Studying this neural network offers several advantages: 1) its convex nature renders it more amenable to analysis compared to non-convex functions, 2) we possess analytical knowledge of the global optimum, and 3) the global optimum exists at a point of non-differentiability—the setup which is particularly relevant to ReLU networks.
+The limitations of applying results like those in Cohen et al.
+(2021) to non-smooth settings highlight the need for new theoretical frameworks.
+In hindsight, it is not particularly surprising that Cohen et al.
+(2021).’s results do not hold even for the differentiable case, as they derive their key result using the assumption of $L-smoothness$ while acknowledging in their discussion that this assumption is not ideal for neural networks.
+We propose using the convex non-differentiable single-layer neural network from Kumar (2023) as a toy problem for theory development.
+This approach offers advantages in terms of analytical tractability and relevance to ReLU networks, potentially leading to insights about optimization around non-differentiable points and informing more robust deep learning optimization techniques.
 6 Conclusion
-In this paper, we highlight how non-differentiability alters training dynamics across three domains: convergence analysis, the LASSO problem, and the Edge of Stability.
-In each case, we show that problems arise because of incorrect assumptions made about the contour of the loss surface.
-Our results suggest the critical need to reach consensus on a set of assumptions which can be made about the loss surface before progress can be made.
-6 Acknowledgements
-We would like to thank Wenyun Zuo, Boris Ginsburg, and several anonymous reviewers for inputs on earlier drafts of this paper.
-We would also like to thank Shripad Tuljapurkar and Marc Feldman for how to best stare at dynamical systems.
-Appendix A Appendix A - Architecture and code for Figure 2
+This paper demonstrates how non-differentiability fundamentally alters neural network training dynamics, challenging widely held assumptions in deep learning optimization.
+Our findings in convergence analysis, LASSO problems, and the Edge of Stability phenomenon reveal that many current practices based on smooth optimization theory can lead to suboptimal or incorrect results in non-smooth settings.
+For the deep learning community, these results underscore the need to reevaluate common techniques in network pruning, learning rate selection, and convergence analysis.
+We call for a collective effort to develop new theoretical frameworks and empirical validations that accurately capture the behavior of non-smooth optimization in deep learning.
+By building on more appropriate foundations, we can potentially unlock more efficient and reliable training algorithms, leading to improved performance across a wide range of applications.
+As the field continues to evolve, it is crucial that both researchers and practitioners approach optimization in neural networks with a more nuanced understanding of non-smoothness and its implications.
+SUPPLEMENTARY MATERIAL
+Appendix A Appendices for section 3
+A.1 Code to generate Figure 2
 The code to construct our plots in Figure 2, was taken from https://github.com/kuangliu/pytorch-cifar/blob/master/models/vgg.py, with three minor modifications.
-First, we removed the batch normalization as part of each layer.
-We did this because, while it is known that batch normalization alters the training dynamics (Ioffe & Szegedy, 2015), and our work is interested in the effect of network smoothness on convergence speed in the absence of confounders.
+First, (in one set of experiments) we removed the batch normalization as part of each layer.
+We did this because, while it is known that batch normalization alters the training dynamics (Ioffe and Szegedy, 2015), and our work is interested in the effect of network smoothness on convergence speed in the absence of confounders.
 Second, we replaced max pool with average pool to ensure differentiability of the layer, and third, we used a simpler layer configuration than one provided in the link.
@@
               64, 64, 64, 64, ’M’],
-Appendix B Proof of proposition 3.1
+A.2 A Geometric Characterization of Strong Duality
+The loss function in 3 exhibits strong duality, which is typically established using Slater’s condition since 0 is an interior point of the domain.
+In this subsection, we provide a geometric characterization of this fact.
+Interestingly, these geometric properties extend beyond our specific case to convex functions with finite global minima.
+We present these general results before specializing to our non-differentiable case.
+A.2.1 Preliminaries and Notation
+Throughout our analysis, we work with extended value convex functions $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{-\infty,\infty\}$.
+This formulation allows us to work directly in $\mathbb{R}^{n}$, as constraints are automatically handled through the extended value representation: points outside the feasible domain simply take infinite values.
+A fundamental object in our study is the epigraph of $f$, defined as $M=\text{Epi}(f)=\{(x,w):f(x)\leq w\}$.
+A key property we utilize is that the epigraph of a convex function is itself a convex set (Section 3.1.7 in Boyd et al.
+(2004)), providing the geometric foundation for our subsequent analysis.
+A.2.2 Geometric Properties of Global Minima
+We show that there exists a non-empty family of non-vertical hyperplanes passing through the point on the epigraph corresponding to the global minimum of $f$.
+Moreover, no other hyperplane containing the epigraph can have a higher intercept on the $(n+1)$st axis.
+To establish this result, we first analyze the case where $f$ has a global minimum at 0.
+The Min Common Point/Max Crossing (MC/MC) framework (for details, see Chapter 4 of Bertsekas (2009)) provides the tools for our analysis:
+  • Min Common point (w*): The point with minimum $(n+1)$th component among all vectors common to $M$ and the $(n+1)$st axis.
+  • Max Crossing (q*): The highest crossing point on the $(n+1)$st axis among all non-vertical hyperplanes containing $M$ in their upper half-space.
+In our specific case, since $f$ has a global minimum at 0, w* also represents the lowest point of the epigraph along the $(n+1)$st axis. Our proofs utilize the following result:
+Proposition A.1.
+Let $M=\text{Epi}(p)$, where $p:\mathbb{R}^{n}\rightarrow[-\infty,\infty]$ is a convex function with $p(0)=w^{*}<\infty$. Then $q^{*}=w^{*}$ if and only if $p$ is lower semicontinuous at 0.
+Proof.
+See corollary to Proposition 4.3.1 in Bertsekas (2009) (Page 148). ∎
+Using this result, we establish strong duality for the case when $f$ has a global minimum at 0:
+Proposition A.2.
+For $f$ as above, $q^{*}=w^{*}=f(0)$.
+Proof.
+Let $g_{k}=f(u_{k})$ for any sequence $u_{k}\rightarrow 0$. Since 0 is the global minimum of $f$, $g_{k}\geq f(0)$. Taking the limit,
+$\liminf_{k\rightarrow\infty}g_{k}\geq f(0)=w^{*},$
+establishing lower semicontinuity at 0. The result follows from Proposition A.1. ∎
+The equality $q^{*}=w^{*}$ suggests special geometric properties at $(0,\ldots,0,w^{*})$. This point, where the epigraph meets the $(n+1)$st axis at its lowest value, warrants closer examination.
+Proposition A.3.
+There exists a non-empty family of non-vertical hyperplanes $\mathcal{H}$ passing through $(0,\ldots,0,f(0))$ that contain $M$ in their upper half-space.
+No other hyperplane containing $M$ in its upper half-space can have a higher intercept on the $(n+1)$st axis.
+Proof.
+We first show that $(0,\ldots,0,w^{*})$ is not an interior point of $M$.
+Suppose, by contradiction, that it is.
+Then there exists $\epsilon>0$ such that $B_{\epsilon}((0,\ldots,0,w^{*}))\subset M$.
+This implies $(0,\ldots,0,w^{*}-\epsilon/2)\in M$, contradicting that $w^{*}$ is the value at the global minimum.
+By the supporting hyperplane theorem, there exists a supporting hyperplane passing through $(0,\ldots,0,w^{*})$, and therefore, $\mathcal{H}$ is non-empty.
+These hyperplanes necessarily intersect the $(n+1)$st axis at $w^{*}$.
+By weak duality (Proposition 4.1.2 in Bertsekas (2009)), $q^{*}\leq w^{*}$, implying these hyperplanes achieve the highest possible intercept.
+To prove non-verticality, suppose by contradiction that some $h\in\mathcal{H}$ is vertical.
+Since $h$ passes through $(0,\ldots,0,w^{*})$, it must have the form $q^{T}\beta=0$ for some $q\in\mathbb{R}^{n}$, where $q\neq 0$.
+Construct $u\in\mathbb{R}^{n}$ by setting $u[i]=\text{sign}(q[i])$ for each component.
+Then $q^{T}u>0$ by construction.
+Note that points $(u,f(u))$ and $(-u,f(-u))$ must belong to $M$, where $f(u)$ and $f(-u)$ represent the height of the epigraph at these points.
+By convexity, at least one of these heights must be finite, while the other may be infinite (but not $\infty$- ∞ as this would violate the existence of a finite global minimum).
+Importantly, the actual values of these heights do not affect our argument.
+Since $M$ must lie in the upper half-space of $h$, we require $q^{T}(-u)\geq 0$.
+However, $q^{T}(-u)=-(q^{T}u)<0$, a contradiction.
+∎
+The geometric intuition from these results is worth noting: at the minimum point, supporting hyperplanes must exist (the epigraph cannot have an interior point there).
+These hyperplanes are constrained in two fundamental ways.
+First, by weak duality, a hyperplane not belonging to $\mathcal{H}$ cannot support $M$ in its upper halfspace and intersect the $(n+1)$st axis above $w^{*}$ any higher intercept would "cut" through the epigraph.
+Second, these hyperplanes cannot be vertical, as this would "slice" the epigraph into two parts, violating the upper half-space requirement.
+A standard result in convex analysis is that all closed proper convex functions show strong duality (see Section 4.2.1 in Bertsekas (2009) for a demonstration of this fact using the MC/MC framework).
+Our conditions are at least as strong as this condition as shown in the following proposition:
+Proposition A.4.
+Any closed proper convex function satisfies the conditions shown above [^9] It is possible that the stronger ”if and only if” condition holds, but we are unable to show the reverse direction, and leave it for future work.
+Proof.
+Since the closedness and lower semicontinuity of $f$ are equivalent (Proposition 1.1.2 in Bertsekas (2009)), the above results follow as long as $w^{*}$ is finite.
+The properness of a closed, convex $f$ can be argued in two ways:
+  1. 1.
+    A closed improper convex function cannot take finite values (Page 11 of Bertsekas (2009)), and therefore, $f$ has to be proper.
+  2. 2.
+    Since $f$ has a finite global minimum, we have $f(x)>-\infty$ for all $x$, and a finite value for at least some $x$. Therefore, $f$ is proper.
+These properties, established for minimum at 0, extend naturally to any point $c\in\mathbb{R}^{n}$ through translation.
+Since translation moves points parallel to coordinate axes without rotation, it preserves all geometric relationships between hyperplanes and the epigraph, particularly the non-verticality property.
+We formalize this in the following proposition:
+Proposition A.5.
+Let $f$ be a convex function with finite global minimum at $c$. Then $q^{*}=w^{*}=f(c)$.
+Proof.
+Through the change of variable $y=x-c$, we obtain $f(x)=f(y+c)=g(y)$, which is convex in $y$ by composition.
+Now, $g$ is a convex function with global minimum at 0, and the results follow from our previous propositions.
+∎
+The connection of our result to Slater’s condition is an interesting question which we leave to future work.
+A.2.3 Geometric Properties at Non-differentiable Global Minima
+The geometric properties discussed above take on striking implications for our loss function (see Equation 3).
+At a differentiable minimum, the epigraph forms a smooth valley, allowing supporting hyperplanes to transition gradually as they pivot around the minimum.
+This smooth structure enables gradients to follow a continuous, predictable path down to the minimum, aligning with classical optimization techniques.
+In stark contrast, when non-differentiability exists at the minimum, the epigraph assumes a distinct, angled structure.
+Here, supporting hyperplanes encounter a geometric break, forcing abrupt transitions rather than a gradual "roll" around the minimum.
+This creates an environment where traditional gradient-based methods falter—they depend on the very smoothness that’s absent here.
+Consider the family of supporting hyperplanes $\mathcal{H}$ at this non-differentiable minimum.
+While hyperplanes still support the epigraph and are never vertical (from our general theory), their orientation shifts suddenly rather than smoothly around the minimum point.
+This discontinuity reveals a fundamental barrier: without a smooth path, gradient-based methods lack the guidance needed to navigate effectively.
+Thus, this geometric characterization highlights why optimization strategies rooted in smooth settings often struggle in non-differentiable landscapes, shedding light on the limitations encountered near such minima.
+Appendix B Appendices for section 4
+B.1 Proof of proposition 4.1
 Proposition 4.1 follows from the two propositions described below.
@@
 The sequence $\{\beta_{t}\}_{t=0}^{\infty}$ described in the previous proposition does not converge. Furthermore, there exists an integer $N_{0}$, such that for every $N\geq N_{0}$, we can write
-$\beta_{n_{k}+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even},\end{cases}$ (13)
+$\beta_{N+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even},\end{cases}$ (13)
 for some $\gamma_{k}\in(0,\alpha\lambda_{1})$ for all $10\leq k\leq P-$ 1.
@@
 The result follows by choosing $N_{0}=\text{max}(n_{0},n_{1}\ldots n_{P-1})$. ∎
-Appendix C NDGM unreliability for the general LASSO problem
+B.2 NDGM unreliability for the general LASSO problem
 In this section, we demonstrate via simulations that the show that the problems demonstrated in our toy LASSO problem hold even for the general setting.
@@
 At the end of the two runs, the optimal solution with $\lambda_{1}=0.1$ has an $L_{1}$ norm of 1.62, and the optimal solution with $\lambda_{1}=10$ has an $L_{1}$ norm of 25.4 i.e., a 1000 fold increase in the value of $\lambda_{1}$ results in a more than 15 fold increase in the $L_{1}$ norm of the optimal solution!
-Appendix D Different NGDM variants and the LASSO
+B.3 Different NGDM variants and the LASSO
 In this section, we demonstrate that various NDGM variants, traditionally thought to converge similarly based on the analysis of differentiable functions, exhibit notably distinct behaviors even when applied to the simple non-differentiable toy LASSO problem outlined in (7).
-A summary of our findings are in Table LABEL:tab:variantcomp.
+A summary of our findings are in Table 1.
 NDGM variant Unique characteristic
@@
 RMSProp Only $\beta_{0}$ influences the eventual value of the sequence. If sequence hits 0, then stays 0.
-D.1 SGD and the LASSO
+Table 1: Unique characteristics of sequences generated on running the different NDGM variants on the LASSO problem described in (7) with an initialization of $\beta_{0}$, and a LASSO penalty of $\lambda_{1}$.
+These sequences are believed to have similar behavior based on the analysis of differentiable loss functions.
+B.3.1 SGD and the LASSO
 Since the gradient of (7) is independent of the batch in question, the results in Proposition B.2 hold for the SGD case as well.
-D.2 NDGM with momentum and the LASSO
+B.3.2 NDGM with momentum and the LASSO
 With the vanilla NDGM, (11) gives the guarantee that once a parameter value reaches zero during an iteration, it remains zero in all future iterations.
 We show that NDGM with momentum does not provide this guarantee.
-Proposition D.1.
+Proposition B.3.
 Let $\{\beta_{t}\}_{t=0}^{\infty}$ denote the sequence generated by running NDGM with momentum on (7), with a constant learning rate, $\alpha$, and the entries in $\beta_{0}$ being uniformly sampled from $[-1,1]$.
@@
 hence the result. ∎
-D.3 RMSProp and the LASSO
+B.3.3 RMSProp and the LASSO
 Proposition B.2 shows that the vanilla NDGM sequence eventually bounces between two vectors which depend on the strength of regularization, $\lambda_{1}$, and the weight initialization, $\beta_{0}$.
@@
 which we use as a base in our analysis. We now provide the key results for RMSProp:
-Proposition D.2.
+Proposition B.4.
 Let $\{\beta_{t}\}_{t=0}^{\infty}$ denote the sequence generated by running the RMSProp recursion described in (21), with a constant learning rate, $\alpha$, and the entries in $\beta_{0}$ being uniformly sampled from $[-1,1]$.
@@
 Since the initializations are uncountable, the probability of the event has measure 0.
-Proposition D.3.
+Proposition B.5.
 The RMSProp sequence described in (21) eventually behaves like a vanilla NGDM sequence with a LASSO penalty of 1 irrespective of the value of $\lambda_{1}$ chosen in the original problem.
@@
 which is the recursion for the vanilla NGDM algorithm with a LASSO penalty of 1 up to first order. ∎
-Appendix E Connections to deep learning training theory:
+B.4 Connections to deep learning training theory:
 The conventional wisdom in deep learning is that different NDGM variants converge at varying rates, but eventually reach similar near-optimal values by the end of training.
@@
 We note that when $\lambda_{1}\ll 1$, the RMSProp training curve lies above the vanilla NDGM training curve, and the reverse is true for $\lambda_{1}\gg 1$.
-Proposition E.1.
+Proposition B.6.
 For the vanilla NDGM sequence we have $0<\liminf_{N\rightarrow\infty}||\beta_{N}||_{1}\leq\limsup_{N\rightarrow\infty}||\beta_{N}||_{1}<P\alpha\lambda_{1}$, where $P$ is the number of rows in $\beta_{0}.$
@@
 […]
-We know from propositions B.2 and D.3 that the vanilla NDGM sequence and the RMSProp sequence eventually bounce (approximately) between two points each.
+We know from propositions B.2 and B.5 that the vanilla NDGM sequence and the RMSProp sequence eventually bounce (approximately) between two points each.
 Let $v_{1}(\beta_{0},\lambda_{1})$ and $v_{2}(\beta_{0},\lambda_{1})$ indicate the points around which the vanilla NDGM sequence bounces, and let $r_{1}(\beta_{0},\lambda_{1})$ and $r_{2}(\beta_{0},\lambda_{1})$ indicate the points around which the RMSProp sequence bounces.
-From propositions E.1 and B.2 we have that $\mbox{min}(\lambda_{1}||v_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||v_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}^{2}$, and $\mbox{max}(\lambda_{1}||v_{1}(\beta_{0},\lambda_{1}||_{1}),\lambda_{1}||v_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}^{2}$.
-From proposition D.3 we have that the RMSProp sequence eventually behaves like a vanilla NDGM sequence with a LASSO penalty of 1, and therefore, we have $\mbox{min}(\lambda_{1}||r_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||r_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}$, and $\mbox{max}(\lambda_{1}||r_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||r_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}$.
+From propositions B.6 and B.2 we have that $\mbox{min}(\lambda_{1}||v_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||v_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}^{2}$, and $\mbox{max}(\lambda_{1}||v_{1}(\beta_{0},\lambda_{1}||_{1}),\lambda_{1}||v_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}^{2}$.
+From proposition B.5 we have that the RMSProp sequence eventually behaves like a vanilla NDGM sequence with a LASSO penalty of 1, and therefore, we have $\mbox{min}(\lambda_{1}||r_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||r_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}$, and $\mbox{max}(\lambda_{1}||r_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||r_{2}(\beta_{0},\lambda_{1})||_{1})<P\alpha\lambda_{1}$.
 Therefore, when $\lambda_{1}\ll 1$, we intuitively expect $\mbox{max}(\lambda_{1}||v_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||v_{2}(\beta_{0},\lambda_{1})||_{1})<\mbox{min}(\lambda_{1}||r_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||r_{2}(\beta_{0},\lambda_{1})||_{1})$, and when $\lambda_{1}\gg 1$ we expect $\mbox{max}(\lambda_{1}||r_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||r_{2}(\beta_{0},\lambda_{1})||_{1})<\mbox{min}(\lambda_{1}||v_{1}(\beta_{0},\lambda_{1}||_{1},\lambda_{1}||v_{2}(\beta_{0},\lambda_{1})||_{1})$, which is what the figure shows.
@@
 Accordingly, when $\lambda_{1}$ is significantly greater than 1, the RMSProp loss is expected to be less than the vanilla NDGM loss in most scenarios at training’s conclusion, as demonstrated in the figure.
-Appendix F Experiments with the LASSO penalty:
-In this section, we provide the graphs and tables resulting from the experiments outlined in Section 3, along with the respective code to generate them.
-F.1 $L_{1}$ norm v/s iteration number using different NDGM variants
-Figure 5: Comparison of the training curves of RMSProp, SGD and Adam for a pretrained VGG16 finetuned using the $L_{1}$ penalized cross entropy loss.
-All 6 training runs are initialized with the same weights, use the same learning rate, and are fed in batches in the same sequence.
-F.2 Percentage of weights less than $10^{-5}$ for different NDGM variants
+Appendix C Experiments with the LASSO penalty:
+In this section, we provide a table describing the percentage of parameters having absolute values less than $10^{-5}$ as a function of the algorithm used, and the strength of regularization.
+We also provide the code to re-create our plots.
+The code is GPT generated and validated for correctness
 Variant $\lambda_{1}=0.002$ $\lambda_{1}=2$
@@
 We note for Adam and SGD that increasing the $L_{1}$ norm can lead to a reduction in the percentage of prunable parameters.
-F.3 Code for generating the above plots (GPT generated and validated for correctness)
 import torch
@@
         l1_norms_curves[name][lambda1] = l1_norms
+Appendix D Appendices for Section 5
+D.1 Proof or Proposition 5.1
+Note that any loss function satisfying (6) has subgradients bounded above by $L$.
+Additionally, as outlined in Section 1 of (Boyd et al., 2003), the deductions derived from the subgradient method hold equivalently for differentiable functions.
+This equivalence arises from the fact that the only admissible value for the subgradient of a differentiable convex function at any given point is, indeed, its gradient.
+Accordingly, using the convergence properties of the subgradient methods with constant step sizes from Section 2 of (Boyd et al., 2003), we have that $\lim_{k\rightarrow\infty}f(x_{k})-f^{*}\leq\alpha L^{2}<\infty$, hence the result.